### v2

# Jet schemes, Newton polygons and continued fractions

Abstract.
We study jet schemes of Newton non-degenerate plane curve singularities. We identify a subgraph of the graph of jet components and show that it can be constructed from walks on the lattice points in the first quadrant of the Cartesian plane. In particular, we determine all the irreducible components of the jet schemes. Furthermore, we prove that this subgraph encodes the embedded topological type of the curve singularity in the plane. Finally, we introduce a generating series defined in terms of the irreducible components of the jet schemes and their (co-)dimensions, and we prove that this series is rational and explicitly determine its poles.
Keywords: jet schemes, curve singularities, newton non-degenerate, generating series
The second author author was partially supported by Fondecyt project 1221535 and Fondecyt Exploración project 13250049.
1. Introduction
The main objects studied in this article are the jet schemes of plane singular curves. For every integer $m\in\mathbf{N},$ the $m$-th jet scheme of a singular variety $X$ (here defined over an algebraically closed field $k$), denoted by $X_{m},$ provides a finite-dimensional approximation of the arc space of $X.$ The arc space, denoted by $X_{\infty},$ is the moduli space parametrizing arcs traced on $X;$ in other words, each point of $X_{\infty}$ corresponds to an arc on $X.$ In a certain sense, the geometry of $X_{m}$ is equivalent (up to a trivial fibration) to the geometry of the family of arcs lying in a smooth ambient space containing $X,$ whose ”order of contact” with $X$ is greater than or equal to $m.$
Jet schemes and arc spaces are known to encode a great deal of information about singularities. Through jet schemes, one can detect rational complete intersection singularities and compute the log canonical threshold of a pair [21, 22]. They also lie at the heart of the mechanism of geometric motivic integration [8], and they allow for the definition of the Igusa motivic zeta function (the motivic analogue of the Igusa zeta function in the $p$-adic setting) [7]. Although their structure is generally rather intricate, the guiding philosophy for studying them can be summarized as follows [18]:
”Arc spaces and jet schemes can transform a difficult problem concerning a relatively simple object into a relatively simple problem concerning a difficult object.”
Their structure has been understood for several classes of singularities, i.e., [14, 15, 9, 17, 4, 20]. A distinctive feature in the study of jet schemes is that, in order to extract information about singularities, one must consider them in a dynamical way: letting $m$ vary and carefully analyzing how the geometry of the jet schemes changes with $m.$ This was one of the motivation of the introduction of the graph of jet components [14, 17]. This is a leveled graph which for a variety $X$ is defined as follows: Its vertices at the level $m$ correspond to the irreducible components of $X_{m}$ and an edge is drawn from a vertex v’ of order $m+1$ to a vertex v’ of order $m$ whenever the component corresponding to $v^{\prime}$ maps into the one corresponding to $v$ under the natural truncation map $X_{m+1}\longrightarrow X_{m}.$ It has been shown, for certain classes of singularities, that the data of this graph (in fact, already of a suitable subgraph) can completely determine the embedded topological type of a singularity [14, 17], and in some cases even its analytical type [4]. Moreover, it is closely related to the problem of resolution of singularities and to the embedded Nash problem [5, 6, 15, 19, 11, 16, 2, 12].
The main purpose of this article is to determine the irreducible components of the jet schemes of a Newton non-degenerate plane curve singularity, and to identify a subgraph of the graph of jet components that encodes the embedded topological type of the singularity. Recall that Newton non-degenerate plane curve singularities are those that can be resolved by a toric morphism whose target is the plane. In the case of an irreducible singularity, we show that this problem is related to the following question: let $p,q$ be two coprime positive integers with $0<p<q.$ Denote by $L_{(p,q)}$ the ray emanating from the origin of the plane in the direction of the point $(p,q),$ i.e., the half-line passing through $(p,q)$ with slope $q/p.$
Let $\alpha\in L_{(p,q)}\cap\mathbf{N}^{2}_{0}$. Determine the walk in the plane obtained by the broken line which starts at the point $\alpha+(1,1)$ and in each step it goes towards the halfline $L_{(p,q)}$ until reaching it if possible.
See Figure 1 for a graphical intuition of the problem. We show that this walk (uniquely determined by the problem) can be described by a special type of continued fraction defined via iterated Euclidean divisions. The subgraph of the graph of jet components that we seek is “almost” precisely given by the walks described above (as $\alpha$ varies), where $p$ and $q$ are determined by the Newton polygon of the (irreducible) curve; the integers $p$ and $q$ are closely related to the Puiseux pairs. In the reducible case, the walk is slightly more involved, but we obtain an analogous result. In particular, we determine the irreducible components of all the jet schemes. We prove that the data of the graph of jet components determine the embedded topological type of the curve.
In the last part, for a plane curve singularity $\mathcal{C}$ as above, we consider the following generating series:
$G=\sum_{K,m}u^{\text{codim}(K;\mathbf{A}^{2}_{\infty})}\cdot v^{m}\in\mathbf{Z}[[u,v]]$
where $m\geq 1$ and $K$ is an irreducible component in $\text{Cont}^{m}(\mathcal{C})^{0}\neq\varnothing.$ Here, $\text{Cont}^{m}(\mathcal{C})^{0}$ is very much related to the (m1)(m-1)-th jet scheme, see section 2 for more details. This series was considered first in the PhD thesis of the last author. We show that $G=G(\mathcal{C})$ is rational in $u,v$ (this means that $G\in\mathbf{Z}(u,v)$) and determine its poles. This series is very much related to other zeta functions associated with singularities [7, 13, 23] and their poles have a lot in common but may be different.
The article is organize as follows:
  • In section 2, we define the jet schemes and arc spaces of a variety over an algebraically closed field $k$. The arc spaces, will in fact, be used in computing the irreducible components of the jet schemes (over the origin) of a Newton non-degenerate plane curve singularity.
  • In section 3, we will describe a staircase algorithm that will determine the irreducible components of these jet schemes.
  • Sections 4 and 5 are dedicated to defining newton non-degenerate branches and curves in the plane $\mathbf{C}^{2}$.
  • In section 6, we prove the results that will be used as lemmas in proving the main theorems in section 7.
  • Finally, in section 8, we compute an invariant: the generating series associated to the jet schemes over the origin of a Newton non-degenerate plane curve singularity and show it is rational and determine its ”poles”.
2. Jet Schemes and Arc Spaces
In this section, we introduce the jet schemes and arc space of a variety.
Denote by $k$ an algebraically closed field. Denote by $\mathbf{Var}_{k}$ the category of varieties over $k$. Let $X\rightarrow Spec(k)$ be such a variety. For $m\in\mathbf{N}_{0}$, denote by $\mathcal{X}_{m}:\mathbf{Var}_{k}\rightarrow\mathbf{Set}$ the functor defined by
$\mathcal{X}_{m}(Z)=\mathbf{Var}_{k}[Z\times_{k}Spec(k[t]/\langle t^{m+1}\rangle),X]$
This functor is representable [24] by a variety $X_{m}$ over $k$ called the jet scheme at the level $m$ of $X$. For $m\geq n$, there is a morphism $\pi_{m,n}:X_{m}\rightarrow X_{n}$ called the truncation morphism that is induced by the natural map of rings
$k[t]/\langle t^{n+1}\rangle\rightarrow k[t]/\langle t^{m+1}\rangle$
These morphisms verify: $\pi_{m,n}\circ\pi_{n,p}=\pi_{m,p}$ for $m\geq n\geq p$. The arc space $X_{\infty}$ of $X$ is the limit object of the jet schemes $X_{m}$ along the truncation morphisms $\pi_{m,n}:X_{m}\rightarrow X_{n}$ for $m\geq n$.
Example 1.
Let’s show how to compute the jet schemes of affine hypersurfaces in $k^{n}=\mathbf{A}^{n}$ using the arc space $\mathbf{A}^{n}_{\infty}$. Set $\mathcal{O}_{k^{n}}=k[\mathbf{x}]$ : $\mathbf{x}=(x_{1},...,x_{n})$ the ring of the affine space $k^{n}$ and let $f\in\mathcal{O}_{k^{n}}$ define a radical ideal. Define $X=\{f=0\}\subset k^{n}$ to be the reduced algebraic set whose ring of functions is
$\mathcal{O}_{X}=\mathcal{O}_{k^{n}}/{\langle f\rangle}$
The ring of the arc space of $k^{n}$ is $\mathcal{O}_{\mathbf{A}^{n}_{\infty}}=k[\mathbf{x}^{(0)},\mathbf{x}^{(1)},\mathbf{x}^{(2)},$...] with variables $\mathbf{x}^{(j)}=(x^{(j)}_{1},...,x^{(j)}_{n})$. Introduce the formal power series:
$x_{i}(t)=\sum_{j=0}^{\infty}x_{i}^{(j)}t^{j}\in\mathcal{O}_{\mathbf{A}^{n}_{\infty}}[[t]]$
Expand $f[x_{1}(t),...,x_{n}(t)]=\sum_{m\geq 0}f^{(m)}t^{m}$. Define now $F_{m}$ to be the algebraic set in $\mathbf{A}^{n}_{\infty}$
$F_{m}=\{f^{(j)}=0\}_{0\leq j\leq m}\subset\mathbf{A}^{n}_{\infty}$
Set $\mathbf{A}^{n}_{m}=\{\mathbf{x}^{(j)}=0\}_{j>m}\subset\mathbf{A}^{n}_{\infty}$. The jet scheme of $X$ at the level $m$ is $X_{m}=\mathbf{A}^{n}_{m}\cap F_{m}\subset\mathbf{A}^{n}_{m}$.
In this study, we take $n=2$, $x=x_{1}$, $y=x_{2}$ and every chosen $f\in\mathcal{O}_{k^{2}}$ is reduced. Set $\mathcal{C}=\{f=0\}$. We are interested in the jet schemes at each level $m$ of $\mathcal{C}$ over the origin of $k^{2}$ ie
$\mathcal{C}^{0}_{m}=\mathbf{A}^{2}_{m}\cap F_{m}\cap\{x^{(0)}=y^{(0)}=0\}\subset\mathbf{A}^{2}_{m}$
That’s why, we will introduce: for $\alpha\in\mathbf{N}^{2}$, $\mathcal{H}^{\alpha}_{\infty}$ to be
$\mathcal{H}^{\alpha}_{\infty}=\{x^{(i)}=y^{(j)}=0:0\leq i<\alpha_{1},0\leq j<\alpha_{2}\}\subset\mathbf{A}^{2}_{\infty}$
Also, set $F^{\alpha}_{m}=\{x^{(\alpha_{1})}\neq 0\}\cap\{y^{(\alpha_{2})}\neq 0\}\cap F_{m}\cap\mathcal{H}^{\alpha}_{\infty}\subset\mathbf{A}^{2}_{\infty}$ and
$\mathcal{H}^{\alpha}_{m}=\mathcal{H}^{\alpha}_{\infty}\cap\mathbf{A}^{2}_{m}\subset\mathbf{A}^{2}_{m}$
Moreover, the contact loci $\text{Cont}^{m}(\mathcal{C})^{0}\subset\mathbf{A}^{2}_{\infty}$ is defined as
$\text{Cont}^{m}(\mathcal{C})^{0}=\{x^{(0)}=y^{(0)}=f^{(0)}=...=f^{(m-1)}=0\}\cap\{f^{(m)}\neq 0\}$
If we embedd $\mathbf{A}^{2}_{m-1}\subset\mathbf{A}^{2}_{\infty}$ naturally for all $m$, then one can write
$\text{Cont}^{m}(\mathcal{C})^{0}=\mathcal{C}^{0}_{m-1}\cap\{f^{(m)}\neq 0\}$
Then, studying the irreducible components of $\mathcal{C}^{0}_{m}$ for every $m$ is equivalent to studying those of $\text{Cont}^{m}(\mathcal{C})^{0}$ for all $m\in\mathbf{N}$.
3. Special Staircase in the Plane
To determine the irreducible components of the jet schemes in the main theorems, a staircase algorithm will be described in this section.
Let $p,q$ be two coprime positive integers such that $0<p<q.$ Let $L_{(p,q)}$ be the halfline which starts at the origin of the plane and whose slope is $q/p;$ it passes through the point $(p,q)$.
We call elementary steps the unit vectors $(1,0)$ (horizontal) and $(0,1)$ (vertical). A broken line is a finite sequence of such elementary steps. In this section we tackle the following question:
Question 1.
Let $\alpha\in L_{(p,q)}\cap\mathbf{N}^{2}_{0}$. What is the walk in the plane obtained by the broken line which starts at the point $\alpha+(1,1)$ and in each step it goes towards the halfline $L_{(p,q)}$ until reaching it if possible ?
Somehow, starting at the point $(1,1)$, the walker is always attracted by the halfline $L_{(p,q)}$ but because of the unit of the steps he is obliged to jump over $L_{(p,q)}$ sometimes before reaching it. Note that the case $0<q<p$ is dual where we replace vertical steps by horizontal ones and vice versa.
Example 2.
Let’s start by observing the example where $p=2$, $q=3$ and $\alpha=(0,0)$: the walk goes along the following points: $(1,1),(1,2),(2,2)$, $(2,3)$ as in Figure 1. This can be encoded by the formula:
$(1,1)+\mathbf{(0,1)+(1,0)+(0,1)}=(2,3)$
[…]
Figure 1. The walk mentioned in question 1 for $p=2$, $q=3$ and $\alpha=(0,0)$
Example 3.
Similarly to the previous example, if we take $p=2$, $q=3$ and $\alpha=(2,3)$, we find the same pattern for $(3,4)=(2,3)+(1,1)$:
$(3,4)+\mathbf{(0,1)+(1,0)+(0,1)}=(4,6)$
To answer Question 1 for $0<p<q$ coprime, we consider $\alpha=(0,0)$ and introduce a form of continued fractions which is given by the following iterated Euclidean divisions :
$q$ $=$ $d_{1}p+r_{1}$
$q+r_{1}$ $=$ $d_{2}p+r_{2}$ (1)
$\vdots$
$q+r_{k-1}$ $=$ $d_{k}p+r_{k}$
$\vdots$
where in the first equality $d_{1}$ and $r_{1}$ are respectively the quotient and the remainder of the Euclidean division of $q$ by $p;$ recursively, for $k\geq 2,$ $d_{k}$ and $r_{k}$ are respectively the quotient and the remainder of the division of $q+r_{k-1}$ by $p;$ in particular we have $0\leq r_{k}<p.$ If we obtain $r_{i}=0$ for some $i\geq 1,$ we stop and we denote the continued fractions
$\mathrm{SC}\left(\frac{q}{p}\right)=[[d_{1},d_{2},\ldots,d_{i}]].$
Here $\mathrm{SC}$ stands for staircase; this name is justified in the theorem below.
Theorem 1.
The walk which answers Question 1 for $\alpha=(0,0)$ is determined by the following steps: starting from $(1,1),$
For $1\leq i\leq p-1$, $d_{i}$ vertical unit steps followed by one horizontal unit step. Finally, $d_{p}-1$ vertical unit steps.
The walk has the form of a staircase and the $d_{k}$’s determine the heights of the steps. This is why we introduced the notation $\mathrm{SC}.$ Before proving the theorem, we need the following lemma.
Lemma 1.
We have the following:
  1. (1)
    For $k=1,\ldots,p-1,$ $r_{k}>0$ and $r_{p}=0;$ in particular
    $\mathrm{SC}\left(\frac{q}{p}\right)=[[d_{1},\ldots,d_{p}]].$
  2. (2)
    $q=d_{1}+\ldots+d_{p}.$
Proof.
Summing the first $k$ equations in (3), and then dividing by $p,$ we get
$\frac{kq}{p}=d_{1}+\cdots+d_{k}+\frac{r_{k}}{p}$ (2)
Since $p$ and $q$ are coprime, for $k<p,$ the left member in equation (2) is not an integer, hence $r_{k}\not=0.$
For $k=p,$ the left side in equation (2) is an integer, hence the right side also is; this implies that $r_{p}=0$ and $q=d_{1}+\ldots+d_{p}.$
[…]
Proof.
We need to prove that the walk goes along the following points
$(1,1),(1,2),\ldots,(1,d_{1}+1),(2,d_{1}+1)$
$(2,d_{1}+1),(2,d_{1}+2),\ldots,(2,d_{1}+d_{2}+1),(3,d_{1}+d_{2}+1)$
$\vdots$
$(p-1,d_{1}+\cdots+d_{p-2}+1),\ldots,(p-1,d_{1}+\cdots+d_{p-1}+1),(p,d_{1}+\cdots+d_{p-1}+1)$
$(p,d_{1}+\cdots+d_{p-1}+2),\ldots,(p,d_{1}+\cdots+d_{p-1}+d_{p}).$
Since we took $p<q,$ the line $L_{p,q}$ is above the point $(1,1),$ and hence the first step should be vertical. It is sufficient to prove that for $k=1,\ldots,p-1,$ we have :
  1. (1)
    The point $(k,d_{1}+\cdots+d_{k})$ is strictly below $L_{p,q}.$
  2. (2)
    The point $(k,d_{1}+\cdots+d_{k}+1)$ is strictly above $L_{p,q}.$
  3. (3)
    The point $(k+1,d_{1}+\cdots+d_{k}+1)$ is strictly below $L_{p,q}.$
  4. (4)
    The point $(p,d_{1}+\cdots+d_{p-1}+d_{p})$ belong to $L_{p,q}$.
The last assertion follows from the equality $q=d_{1}+\ldots+d_{p}$ in lemma 1. Here the expression the point is strictly below (respectively above) $L_{p,q}$ means that the point belongs to the relative interior of the cone generated by $(1,0)^{t}$ and $(p,q)^{t}$ (respectively by $(p,q)^{t}$ and $(0,1)^{t})$. The first assertion is equivalent to $\det((k,d_{1}+\cdots+d_{k})^{t},(p,q)^{t})>0;$ this latter determinant is equal to
$kq-pd_{1}-\cdots-pd_{k}=$
$kq-(q-r_{1})-(q+r_{1}-r_{2})-\cdots-(q+r_{k-1}-r_{k})=r_{k};$
we have simply replaced $pd_{i}$ by $q+r_{i-1}-r_{i},$ see the equations (3). Since by lemma 1 we have $r_{k}>0,$ we obtain the first assertion. Similarly,
$\det((p,q)^{t},(k,d_{1}+\cdots+d_{k}+1)^{t})=p-r_{k}>0$
hence we have the second assertion. The last assertion follows from the fact that $\det((k+1,d_{1}+\cdots+d_{k}+1)^{t},(p,q)^{t})=q-p+r_{k}>0$. ∎
Finally, to answer Question 1 for all $\alpha$, consider the corollary:
Corollary 1.
Let $0<p<q$ be coprime integers and $\alpha\in L_{(p,q)}\cap\mathbf{N}^{2}_{0}$. Then, Question 1 for $(p,q,\alpha)$ and $(p,q,\alpha_{0})$ : $\alpha_{0}=(0,0)$ have the the same sequence of elementary steps.
Proof.
Write $\alpha=a\cdot(p,q)$ with $a\geq 0$ such that $\alpha+(1,1)=(ap+1,aq+1)$.
$(p,q)=(1,1)+\sum_{i=1}^{p-1}[d_{i}\cdot(0,1)+(1,0)]+(d_{p}-1)(0,1)$
Also, since $\alpha\in L_{(p,q)}$ then $\alpha+(p,q)\in L_{(p,q)}$ and we obtain:
$\alpha+(1,1)+\sum_{i=1}^{p-1}[d_{i}\cdot(0,1)+(1,0)]+(d_{p}-1)(0,1)=\alpha+(p,q)$
Which concludes the proof. ∎
4. Jet Schemes of Newton Plane Branches
In this section, we are interested in jet schemes of Newton non-degenerate plane branches. We begin by introducing combinatorial objects which are very important in this article: Newton polygons. A plane curve singularity $\mathcal{C}$ is the zero locus of an (reduced) element
$f=\sum c_{\alpha\beta}x^{\alpha}y^{\beta}\in\mathcal{O}_{k^{2}}$
where $\mathcal{O}_{k^{2}}=k[x,y]$ has variables $x=x_{1}$ and $y=x_{2}$. One can associate with $\mathcal{C},$ more precisely with any $f\in\mathcal{O}_{k^{2}}$ its Newton polygon at the origin (notice that $f(0)=0$) as follows; let
$\Gamma^{+}(f,x,y)=\mbox{Conv}\{(\alpha,\beta)+\mathbf{R}_{\geq 0}^{2};\>\ c_{\alpha\beta}\not=0\}\subset\mathbf{R}_{\geq 0}^{2}$
where Conv stands for the convex hull. The Newton polygon $\Gamma(f,x,y)$ is the union of the compact faces of (the boundary) of $\Gamma^{+}(f,x,y);$ it has one dimensional faces and vertices (00-dimensional faces). The dual fan of $\Gamma(f,x,y)$ (the dual Newton fan) is a fan (a collection of cones) whose support is $\mathbf{R}_{\geq 0}^{2}$ ($\mathbf{R}^{2}$ is the dual of the $\mathbf{R}^{2}$ that contains $\Gamma^{+}(f,x,y))$ and which is defined as follows: take $\omega\in\mathbf{R}^{2}_{\geq 0};$ and let
$\nu_{\omega}(f)=\min\{\omega(\alpha,\beta);~~c_{\alpha\beta}\not=0\},$
where $\omega(\alpha,\beta)$ can be thought as the scalar product of $\omega$ and $(\alpha,\beta).$ The supporting hyperplane associated with $\omega$ is
$H_{\omega}(f)=\{(\alpha,\beta)\in\mathbf{R}^{2};\omega(\alpha,\beta)=\nu_{\omega}(f)\}$
With a face $\Phi$ of NP(f,x,y),(f,x,y), one associates the cone
$C_{\Phi}:=\{\omega\in\mathbf{R}^{2};\Phi\subset H_{\omega}(f)\};$
here $\Phi$ can be zero dimensional. The set $\{C_{\Phi},\Phi~\mbox{a face of NP}(f,x,y)\}$ is the dual Newton fan. The tropical variety $T(f)$ of $f$ is the union of the cones $C_{\Phi}$ associated to a compact segment $\Phi$ of the Newton Polygon of $f$. Let $p$ and $q$ be as in Section 3. A newton non-degenerate plane branch (or simply a Newton plane branch) $\mathcal{C}$ is the zero locus of an element
$f=y^{p}-cx^{q}+\sum_{pa+qb>pq}c_{ab}x^{a}y^{b}\in\mathcal{O}_{k^{2}}$
This coincides with the definition which we will see later in Section 5. In the case of $f$ defining a plane Newton non-degenerate branch, NP(f,x,y)(f,x,y) has two vertices $(q,0)$ and $(0,p)$ and one compact face which is the segment joining these two vertices. The dual of $(q,0)$ (respectively $(0,p))$ is the two dimensional cone generated by $(p,q)$ and $(0,1)$ (respectively $(1,0)$ and $(p,q))$. The dual cone of the compact segment is the one dimensional cone generated by $(p,q)$. This 1-dimensional cone is in fact $L_{(p,q)}$.
Set $\tau_{m+1}=\pi_{m+1,m}:\mathcal{C}_{m+1}\rightarrow\mathcal{C}_{m}$ the truncation morphism for $m\in\mathbf{N}_{0}$. Denote $irr(\mathcal{C}^{0}_{m})$ the set of irreducible components of the reduced scheme $\mathcal{C}^{0}_{m}$. Define the jet-component graph $Jet(f)$ of $f$ as follow:
  • the vertices of $Jet(f)$ are the elements of $\bigcup_{m\in\mathbf{N}_{0}}irr(\mathcal{C}^{0}_{m})$,
  • each vertex is weighted by $(d,e)$ where $d$ is its Krull dimension and $e$ its embedding dimension,
  • an edge of $Jet(f)$ joins two irreducible components $V_{m}\in irr(\mathcal{C}_{m}^{0})$ and $V_{m+1}\in irr(\mathcal{C}_{m+1}^{0})$ for some $m\in\mathbf{N}_{0}$, such that we have:
    $\tau_{m+1}(V_{m+1})\subseteq V_{m}$
An element of $irr(\mathcal{C}^{0}_{m})$ is called vertex of $Jet(f)$ at the level $m\in\mathbf{N}_{0}$.
Example 4.
Let’s draw the jet components graph of the cusp ie $f=y^{2}-x^{3}$ in $\mathbf{A}^{2}$ (Figure 2) in two different ways. In the drawing on the right, introduce the notation $(\alpha,w)=\mathcal{H}^{\alpha}_{w}$. Note that we can have $w=\nu_{\alpha}(f)-1$ or $w=\infty$.
225566778811mm(2,2)(2,2)(7,7)(7,7)(8,9)(8,9)(3,3)(3,3)(9,9)(9,9)(9,11)(9,11)(10,13)(10,13)$(10,10)$
[…]
Figure 2. Different representations of the jet components graph of the cusp.
The drawing on the left is obtained by the way that is described in the definition in the previous page. The one on the right is the new representation of the jet components graph that we’re introducing in this article. What makes the drawing on the right better than the one on the left are the following ideas:
  • The drawing on the right is obtained by an algorithm whose steps are well-known. Moreover, it describes a link between the tropical variety and the irreducible components of the jet schemes (which isn’t the case for the drawing on the left).
  • The algorithm mentioned in the previous point is the same for any Newton non-degenerate plane branch (which we will see in the next example).
  • When $k=\mathbf{C}$, the graph on the left is equivalent data with the embedded topological type of the singularity (defined by the cusp) as proven in [14]. Whereas the graph on the right gives richer algebraic data such as an exact description of the $\alpha$ (appearing in the irreducible components of the form $\mathcal{H}^{\alpha}_{m}$).
  • We have to add to the graph on the left weights $(d,e)$ for each vertex in order to know which component vanishes at a certain level (that is when $d=e$) and which doesn’t. However, in the graph on the right, all nonvanishing components lie on the tropical variety.
  • We can obtain the graph on the left from the one on the right: for every edge $\alpha-\alpha^{\prime}$ appearing in the graph on the right; where $w=\nu_{\alpha}(f)-1$ and $w^{\prime}=\nu_{\alpha^{\prime}}(f)-1$ are the corresponding weights $w<w^{\prime}$, add on the edge $\alpha-\alpha^{\prime}$ an amount of $w^{\prime}-w-1$ dots corresponding to the $\mathcal{H}^{\alpha}_{m}$ where $w\leq m<w^{\prime}$.
According to [14], every irreducible component $V_{m}\subset\mathcal{C}^{0}_{m}$ can be written as $V_{m}=\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha}_{m}}$ for a unique $\alpha\in\mathbf{N}^{2}$. In fact, using the notations in [14], if $\nu_{\alpha}(f)>m$ then $V_{m}=B_{m}$ and if $\alpha\in L_{(2,3)}$, $\nu_{\alpha}(f)\leq m$ then $V_{m}=\overline{Cont^{\alpha_{1}}(x)_{m}}$ doesn’t vanish and continues till $\infty$. This comes from: $\nu_{\alpha}(f)=6r$ when $\alpha=(2r,3r)$ : $r\in\mathbf{N}$.
In the representation, we can simply replace $\mathcal{H}^{\alpha}_{m}$ by $\alpha$ and remove the text in red (for $f=y^{2}-x^{3}$) that is the name of an irreducible component that doesn’t vanish. This is because there is a correspondance between the irreducible components of $\mathcal{C}^{0}_{m}$ : $m\in\mathbf{N}$ and some weighted points in $\mathbf{R}_{\geq 0}^{2}$. Here is how it is defined: let $V_{m}\subset\mathcal{C}^{0}_{m}$ be an irreducible component written as $V_{m}=\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha}_{m}}$ for a unique $\alpha\in\mathbf{N}^{2}$. We associate to $V_{m}$ the weighted point $(\alpha,\infty)$ if $\alpha\in L_{(2,3)}$, $\nu_{\alpha}(f)\leq m$ and otherwise, we associate to $V_{m}$ the weighted point $\alpha,\nu_{\alpha}(f)-1]$.
Irreducible components Interpretation Jet-components graph
$\mathcal{H}^{\alpha}_{m}:m<\nu_{\alpha}(f)$ Big, Vanishing $\alpha,\nu_{\alpha}(f)-1]$
$\mathbf{A}^{2}_{m}\cap F^{\alpha}_{m}:\alpha\in T(f)$ Nonvanishing $(\alpha,\infty)$
Table 1. Dictionary of notations for $f=y^{2}-x^{3}$.
The weight of the point $\alpha$ is the level $w$ of the jet scheme $\mathcal{C}^{0}_{w+1}$ where the irreducible component $V_{w+1}$ vanishes: if $w<\infty$, $V_{w}=\mathcal{H}^{\alpha}_{w}$ is replaced by either $\mathcal{H}^{\alpha+(0,1)}_{w+1}$ [when $\alpha$ is below $L_{(2,3)}$] or $\mathcal{H}^{\alpha+(1,0)}_{w+1}$ [when $\alpha$ is above $L_{(2,3)}$]. If $w=\infty$, then $V_{m}$ doesn’t vanish.
We notice that if we unweight the points $(\alpha,w)\mapsto\alpha$ then these $\alpha$ follow the same pattern as in Figure 1. This concerns $f=y^{2}-x^{3}$.
Example 5.
Let’s move on to any $f$ defining a Newton non-degenerate plane branch with $L_{(p,q)}$ as a tropical variety and $0<p<q$ coprime: let $\lambda\in L_{(p,q)}\cap\mathbf{N}^{2}_{0}$. Figure 3 is the staircase representation $J_{SC}(f)$ associated to $f$.
[…]
Figure 3. Drawing of the unweighted points representing the irreducible components of $\mathcal{C}^{0}_{m}$ : $m\in\mathbf{N}$. The red arrow corresponds to an irreducible component that continues to infinity.
Here is a description of $J_{SC}(f)$ in Figure 3: let $\lambda\in L_{(p,q)}\cap\mathbf{N}^{2}_{0}$ and recall the staircase notation $d_{1},...,d_{p}]]$ associated to $(p,q)$ from Theorem 1:
  • Start at $\lambda+(1,1)$;
  • For $1\leq i\leq p-1$, move $d_{i}$ vertical unit steps followed by one horizontal step;
  • Finally, move $d_{p}-1$ vertical steps; we reach $\lambda+(p,q)$;
  • Add a red arrow above $\lambda+(p,q)$;
  • Start over from $\lambda+(p,q)+(1,1)$.
Note that each time we move from a point $\alpha$ to $\alpha+(0,1)$ or $\alpha+(1,0)$, we delete $\alpha$ and we replace it by the new point. We weight each point $\alpha$ by $\nu_{\alpha}(f)-1$. The red arrows are never deleted and they are associated to points on the tropical variety with weight $\infty$. In the case where $0<q<p$, vertical steps are replaced by horizontal steps and vice versa.
The below proposition shows that $J_{SC}(f)$ determines $Jet(f)$.
Proposition 1.
If $f,g$ define Newton plane branches with $J_{SC}(f)=J_{SC}(g)$ then we have $Jet(f)=Jet(g)$.
Proof.
Consider first an edge in Figure 3. Denote by $(\alpha,w)$ and $(\alpha^{\prime},w^{\prime})$ with $w<w^{\prime}<\infty$ the weighted vertices of this edge. Note that $\alpha^{\prime}=\alpha+(1,0)$ or $\alpha^{\prime}=\alpha+(0,1)$ with $w=\nu_{\alpha}(f)-1$ and $w^{\prime}=\nu_{\alpha^{\prime}}(f)-1$. Add on the edge $\alpha\alpha^{\prime}$, $w^{\prime}-w-1$ vertices representing the big components defined by the same $\alpha$.
Now consider a nonvanishing component represented by $(\alpha,\infty)$. Add on the arrow above $\alpha$ a countably infinite amount of vertices. These represent the nonvanishing components for $m\geq\nu_{\alpha}(f)$ and $\alpha\in T(f)$. Finally, scale each weight $w$ by its level $m=w+1$.
The construction we just defined allows to obtain $Jet(h)$ from $J_{SC}(h)$ for every polynomial (or power series) $h$ which is non-degenerate (with respect to $x$ and $y$). Also, the construction doesn’t really depend on $h$ but on the weighted graph $J_{SC}(h)$. This allows to conclude that since $J_{SC}(f)=J_{SC}(g)$ then $Jet(f)=Jet(g)$ . ∎
5. Jet Schemes of Newton Plane Curves
We call $\mathcal{C}\subset k^{2}$ a Newton non-degenerate plane curve (or simply a Newton plane curve) if it is the zero locus of some reduced
$f=\sum_{(\alpha,\beta)\in supp(f)}c_{\alpha,\beta}\cdot x^{\alpha}y^{\beta}\in\mathcal{O}_{k^{2}}$
such that for every face $\Phi$ of the Newton Polygon of $f$, if we set
$f_{\Phi}=\sum_{(\alpha,\beta)\in\Phi\cap supp(f)}c_{\alpha,\beta}\cdot x^{\alpha}y^{\beta}$
then the system of equations:
$f_{\Phi}=\frac{\partial f_{\Phi}}{\partial x}=\frac{\partial f_{\Phi}}{\partial y}=0$
has no solutions in the torus $(k^{\times})^{2}$.
In this study, it is crucial to know the description of Newton plane curves in order to compute their jet schemes over the origin of $k^{2}$. In this section, we will describe the equations of these curves.
Proposition 2 describes the equation of a Newton plane branch up to an isomorphism.
Proposition 2.
[10] The Newton plane branches are up to an isomorphism the plane curves defined by
$f=y^{p}-cx^{q}+\sum_{pa+qb>pq}c_{a,b}x^{a}y^{b}\in\mathcal{O}_{k^{2}}$
with $p,q\geq 0$ coprime integers (and $f$ has only one Puiseux characteristic pair).
Every curve contained in a Newton plane curve is also Newton non-degenerate. So, now, we describe the equations of Newton Plane curves through their defining polynomials (or power series) as well as the (equational) relations between their different branches.
Theorem 2.
Newton Plane curves $\mathcal{C}\subset k^{2}$ are defined (up to an isomorphism) as such:
  1. (1)
    each branch $\mathcal{B}\subset\mathcal{C}$ is defined by a $g\in\mathcal{O}_{k^{2}}$ such that
    $g=y^{p}-cx^{q}+\sum_{pa+qb>pq}c_{a,b}x^{a}y^{b}$
    as in theorem 2.
  2. (2)
    Let $g^{\prime}=y^{p^{\prime}}-c^{\prime}x^{q^{\prime}}+\sum_{p^{\prime}a+q^{\prime}b>p^{\prime}q^{\prime}}c^{\prime}_{a,b}x^{a}y^{b}$ define another branch of $\mathcal{C}$. If $(p,q)=(p^{\prime},q^{\prime})$ then $c\neq c^{\prime}$.
In order to prove Theorem 2, we refer from [1] to an equivalent definition for Newton non-degenerate polynomials which is the following: let $f=\sum_{a,b}f_{a,b}\cdot x^{a}y^{b}\in k[x,y]$ be reduced and $w\in\mathbf{R}_{\geq 0}^{2}$, the initial form
$in_{w}(f)=\sum_{(a,b)\cdot w=\nu_{w}(f)}f_{a,b}\cdot x^{a}y^{b}$
where $(a,b)\cdot w=aw_{1}+bw_{2}$ is the natural inner product. Then, $f$ defines a Newton Plane Curve if and only if for all $w\in\mathbf{R}_{\geq 0}^{2}$, the reduced algebraic set defined by $in_{w}(f)$ has no singularities in the torus $(k^{\times})^{2}$.
Equivalently, $f$ is Newton non-degenerate if and only if for all $(p,q)\in T(f)\cap\mathbf{N}^{2}$ coprime, $in_{(p,q)}(f)$ is the product of a non-zero constant by a monomial in $x,y$ (which could be equal to $1$) and a finite product of binomials in $x,y$: $\prod_{j}(y^{p}-c_{j}x^{q})$ where $j\mapsto c_{j}$ is injective.
Now, that we know the equations of the branches of a Newton plane curve and how they are related, let’s look at the irreducible components of $\mathcal{C}^{0}_{m}$ when $f=(y^{2}-x^{3})(y^{3}-x^{2})$ as seen in Figure 4. The tropical variety of $f$ is $T(f)=L_{(2,3)}\cup L_{(3,2)}$. Denote by $\sigma=L_{(2,3)}+L_{(3,2)}$ the cone such that $\partial\sigma=T(f)$.
Example 6.
The special staircase walk for $f=(y^{2}-x^{3})(y^{3}-x^{2})$ determines the irreducible components of $\mathcal{C}^{0}_{m}$ : $m\in\mathbf{N}$.
[…]
Figure 4. Special Staircase for $f=(y^{2}-x^{3})(y^{3}-x^{2})$
In fact, every irreducible component $V_{m}\subset\mathcal{C}^{0}_{m}$ is written as $V_{m}=\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha}_{m}}$ for a unique $\alpha\in\mathbf{N}^{2}$. We associate to $V_{m}$ then the weighted point $(\alpha,\infty)$ if $\alpha\in T(f)$, $\nu_{\alpha}(f)\leq m$ and otherwise, we associate to $V_{m}$ the weighted point $\alpha,\nu_{\alpha}(f)-1]$. Thus, we obtain Figure 4 by obeying to the following rules:
  • Start always at $(1,1)$;
  • If $\alpha\in\overset{\circ}{\sigma}$, remove $\alpha$ to replace it with $\alpha+(0,1)$ and $\alpha+(1,0)$;
  • If $\alpha\in L_{(2,3)}$, add a red arrow above $\alpha$ and add $\alpha+(1,0)$;
  • If $\alpha$ is above $L_{(2,3)}$ replace $\alpha$ by $\alpha+(1,0)$;
  • If $\alpha\in L_{(3,2)}$, add a red arrow above $\alpha$ and add $\alpha+(0,1)$;
  • If $\alpha$ is below $L_{(3,2)}$ replace $\alpha$ by $\alpha+(0,1)$.
The weighted point $(\alpha,w)$ represents the irreducible components $V_{m}=\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha}_{m}}$ if $m<w<\infty$ that ”vanish” at the level $w+1$ and if $w=\infty$, then $V_{m}$ continue till $w=\infty$ without vanishing. These $V_{m}$, that do not vanish, are represented by a red arrow.
Now, let’s define the order of a power series $h=\sum_{k}h_{k}t^{k}\in k[[t]]\backslash\{0\}$. We set $ord_{t}(h)$ to be the least exponent $e$ such that $h_{e}\neq 0$. Otherwise if $h=0$, we set
$ord_{t}(0)=\infty$
Recall, now, from [3] the definition of the semigroup of a reduced plane curve with $r$ branches. Let $f=f_{1}...f_{r}$ be a reduced polynomial defining such a curve with $k=\mathbf{C}$. Choose $\phi_{i}:\mathbf{C}_{t}\rightarrow\mathbf{C}_{x,y}^{2}$ a parametrisation (with variable $t$) of $\{f_{i}(x,y)=0\}$, obtained by the Newton-Puiseux theorem. Denote by $v_{i}(g)=(f_{i},g)_{0}=ord_{t}(g\circ\phi_{i})$ for $g\in\mathbf{C}[[x,y]]$.
The semigroup of $f$ is the subsemigroup $\mathcal{S}(f)\subseteq\mathbf{N}^{r}_{0}$ whose elements are the:
$v(g)=[v_{1}(g),...,v_{r}(g)]$
with $v_{i}(g)<\infty$ for all $g\in\mathbf{C}[x,y]$ and index $i$.
It is known [25] that the semigroup determines the embedded topological type of the singularity defined by $f$. When $f$ is Newton non-degenerate, we can rewrite this fact as:
Theorem 3.
Let $f=f_{1}...f_{r}\in\mathcal{O}_{k^{2}}$ be Newton non-degenerate product of $r$ branches and let $g=g_{1}...g_{r}$ be another such polynomial with the same number of branches. Assume
$T(f_{i})=T(g_{i})$
for $1\leq i\leq r$ and for $j\neq i$,
$(f_{i},f_{j})_{0}=(g_{i},g_{j})_{0}$
are equal. Then $f$ and $g$ have the same embedded topological type of the singularity at 0.
6. Useful Results
In this part of the article, we will mention the key results that will be useful in proving the theorems in the next sections.
For $\alpha\in\mathbf{R}^{2}$, set $|\alpha|=\alpha_{1}+\alpha_{2}$.
Moreover, we say that two polynomials $f,g\in k[x_{1},...,x_{n}]$ are congruent modulo a variety $W\subset\mathbf{A}^{\infty}$ (which we write $f=g\mod W$) if they are congruent modulo the defining ideal of $W$ in $k[x_{1},x_{2},$...].
This allows us to introduce the terminology, $f=g\mod\alpha$ when $W=\mathcal{H}^{\alpha}_{\infty}$. Let’s start with this observation which will be used later.
Observation 1.
Let $f\in\mathcal{O}_{k^{2}}$ define a non-degenerate Newton plane branch. Let $\alpha=s(p,q)\in T(f)$ with $gcd(p,q)=1$. Then, $\nu_{\alpha}(f)>|\alpha|$ if $p,q>1$.
Proof.
Write $in_{\alpha}(f)=y^{p}-cx^{q}$. Then,
$\nu_{\alpha}(f)=spq>s(p+q)=|\alpha|$
This proves our observation. ∎
The below lemma will be used (in the next section) in proving Theorem 7.
Lemma 2.
Let $f\in\mathcal{O}_{k^{2}}$ define a non-degenerate plane curve with $T(f)\simeq\mathbf{R}_{\geq 0}$. Set $f=f_{1}...f_{r}$ to be the decomposition of $f$ into branches. Let $\alpha\in T(f)\cap\mathbf{N}^{2}$ and $m\geq\nu_{\alpha}(f)$. Then we have
$F^{\alpha}_{m}=\bigsqcup_{i=1}^{r}F^{\alpha}_{i,\nu_{\alpha}(f_{i})+m-\nu_{\alpha}(f)}$
Proof.
By induction on $m\geq\nu_{\alpha}(f)$.
  • At $m=\nu_{\alpha}(f)$, we have the following:
    $f^{[\nu_{\alpha}(f)]}=f_{1}^{[\nu_{\alpha}(f_{1})]}...f_{r}^{[\nu_{\alpha}(f_{r})]}\mod\alpha$
    and since $T(f)$ is the semi-line that passes by $(0,0)$ and $\alpha$ then for all $i$:
    $f_{i}^{[\nu_{\alpha}(f_{i})]}=in_{\alpha}(f_{i})(x^{(\alpha_{1})},y^{\alpha_{2}})\mod\alpha$
    which isn’t a monomial. Moreover, $in_{\alpha}(f_{j})$ are pairwise coprime as polynomials since $f$ is Newton non-degenerate. This allows to say that:
    $F^{\alpha}_{\nu_{\alpha}(f)}=\bigsqcup_{i=1}^{r}F^{\alpha}_{i,\nu_{\alpha}(f_{i})}$
    This proves the initial case $m=\nu_{\alpha}(f)$.
  • Suppose that for m1m-1, we have:
    $F^{\alpha}_{m-1}=\bigsqcup_{i=1}^{r}F^{\alpha}_{i,\nu_{\alpha}(f_{i})+m-1-\nu_{\alpha}(f)}$
    Then, as $F^{\alpha}_{m}=F^{\alpha}_{m-1}\cap\{f^{(m)}=0\}$, let’s show that:
    $F^{\alpha}_{i,\nu_{\alpha}(f_{i})+m-1-\nu_{\alpha}(f)}\cap\{f^{(m)}=0\}=F^{\alpha}_{i,\nu_{\alpha}(f_{i})+m-\nu_{\alpha}(f)}$
    This comes from the fact that if we set $k=\nu_{\alpha}(f_{i})+m-\nu_{\alpha}(f)$ then
    $f^{(m)}=f^{(k)}_{i}\prod_{j\neq i}f^{[\nu_{\alpha}(f_{i})]}_{i}\mod F^{\alpha}_{i,k-1}$
    So the proposition is true for $m$.
This concludes the induction. ∎
We will show (in the next section) that the irreducible components of the jet schemes $\mathcal{C}^{0}_{m}$ (over the origin) of Newton plane curves $\mathcal{C}=\{f=0\}\subset\mathbf{A}^{2}$ are of two kinds:
  • the ones of the form $\mathcal{H}^{\alpha}_{m}$ who are irreducible and have $|\alpha|$ as codimension. We take the $\alpha$ such that there is no $\beta\leq\alpha$ component-wise with $\alpha\neq\beta$. These $\mathcal{H}^{\alpha}_{m}$ are called having hyperplane coordinates;
  • the ones of the form $\overline{G^{\alpha}_{k}\cap\mathbf{A}^{2}_{m}}$ where $\{g=0\}\subset\mathcal{C}$ is a branch, $m\geq\nu_{\alpha}(f)$, $k=\nu_{\alpha}(g)+m-\nu_{\alpha}(f)$ and $\alpha\in T(g)\cap\mathbf{N}^{2}$; these are called infinite components.
The next lemma will allow us to have their codimensions and verify their irreducibility.
Lemma 3.
Let $f\in\mathcal{O}_{k^{2}}$ define a Newton plane branch. Let $\alpha\in T(f)\cap\mathbf{N}^{2}$ and $\nu_{\alpha}(f)\leq k\leq m$. Then, we have
$codim(F^{\alpha}_{k}\cap\mathbf{A}^{2}_{m};\mathbf{A}^{2}_{m})=|\alpha|+k-\nu_{\alpha}(f)+1$
and $F^{\alpha}_{k}\cap\mathbf{A}^{2}_{m}$ is irreducible.
Proof.
Denote by $\pi:\mathbf{A}^{2}_{m}\rightarrow\mathbf{A}^{2}_{k}$ the truncation morphism of the jet schemes of $\mathbf{A}^{2}$ : they are defined as $\pi(x_{0},...,x_{m},y_{0},...,y_{m})=(x_{0},...,x_{k},y_{0},...,y_{k})$ where $v_{i}=v^{(i)}$ if $v=x$ or $v=y$. Notice that:
$F^{\alpha}_{k}\cap\mathbf{A}^{2}_{m}=\pi^{-1}(F^{\alpha}_{k}\cap\mathbf{A}^{2}_{k})$
So to have $F^{\alpha}_{k}\cap\mathbf{A}^{2}_{m}$ irreducible, it’s enough to have $F^{\alpha}_{k}\cap\mathbf{A}^{2}_{k}$ irreducible. Since $f$ defines a Newton plane branch, then $F^{\alpha}_{k}\cap\mathbf{A}^{2}_{k}$ is irreducible by Lemma 4.3 of [14] and using our notations it has codimension $|\alpha|+k-\nu_{\alpha}(f)+1$. If we lift to level $m$ the irreducibility is preserved as well as the codimension. ∎
Now that we know who are the candidates to be irreducible components of $\mathcal{C}^{0}_{m}$ for all $m$, we need to compare (by inclusion) these candidates to be certain that they have the property of being components. Their irreducibility has been proven in the previous result.
First, let’s recall some notations: let $\mathbf{A}^{2}_{m}$ be the $m$-jet space of $\mathbf{A}^{2}$ with variables $x^{(0)},...,x^{(m)},y^{(0)},...,y^{(m)}$. Let $\alpha\in\mathbf{N}^{2}$.
  • Set $\mathcal{H}^{\alpha}_{m}=\{x^{(i)}=y^{(j)}=0:0\leq i<\alpha_{1},0\leq j<\alpha_{2}\}\subset\mathbf{A}^{2}_{m}$;
  • Let $g$ define a branch in $\mathbf{A}^{2}$ and $\nu_{\alpha}(g)\leq k\leq m$, set
    $\mathbf{A}^{2}_{m}\cap G^{\alpha}_{k}=\{x^{(\alpha_{1})}\neq 0\}\cap\{y^{(\alpha_{2})}\neq 0\}\cap\mathcal{H}^{\alpha}_{m}\cap\{g^{[\nu_{\alpha}(g)]}=...=g^{(k)}=0\}$
Let’s sum up the purpose of this section by Table 2: whenever $\mathcal{H}^{\gamma}_{m}$ is mentioned we impose the condition $\nu_{\gamma}(f)>m$ and whenever
$\mathbf{A}^{2}_{m}\cap F^{\alpha}_{i,a}$
corresponds to a branch defined by a factor $f_{i}$ of $f$ with $\alpha\in T(f_{i})$, we impose $m\geq\nu_{\alpha}(f)$ and $a=\nu_{\alpha}(f_{i})+m-\nu_{\alpha}(f)$.
AA $B$ when does $A\subseteq B?$ why?
$\mathcal{H}^{\alpha}_{m}$ $\mathcal{H}^{\beta}_{m}$ $\alpha\geq\beta$ component-wise
$\mathcal{H}^{\gamma}_{m}$ $\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha}_{i,a}}$ $\gamma=\alpha+(1,1)$, $m=\nu_{\alpha}(f)$
$\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha}_{i,a}}$ $\mathcal{H}^{\gamma}_{m}$ never otherwise, $\alpha\geq\gamma$ absurd
$\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha}_{i,a}}$ $\overline{\mathbf{A}^{2}_{m}\cap F^{\beta}_{i,b}}$ $\alpha=\beta$ and $f_{i}=f_{j}$ compare codimensions
Table 2. Table describing when the candidates $A,B\subseteq\mathcal{C}^{0}_{m}$ verify $A\subseteq B$.
Theorem 4.
Let $f\in\mathcal{O}_{k^{2}}$ define a Newton plane curve and $f_{i}$ define branch of $f$. Write $T(f_{i})=\mathbf{R}_{\geq 0}(p,q)$. Let $\alpha,\beta\in T(f_{i})\cap\mathbf{N}^{2}$ and set $\gamma=\beta+(1,1)$ and suppose $\nu_{\gamma}(f)>m\geq\nu_{\alpha}(f)$. Then, we have
$\mathcal{H}^{\gamma}_{m}\subseteq\overline{F^{\alpha}_{i,\nu_{\alpha}(f_{i})+m-\nu_{\alpha}(f)}\cap\mathbf{A}^{2}_{m}}$
if and only if $\alpha=\beta$ and $m=\nu_{\alpha}(f)$. Moreovoer, this inclusion is always strict.
Proof.
Suppose first $\alpha=\beta$ and that: $\mathcal{H}^{\gamma}_{m}=\overline{F^{\alpha}_{i,\nu_{\alpha}(f_{i})+m-\nu_{\alpha}(f)}\cap\mathbf{A}^{2}_{m}}$. So,
$U:=F^{\alpha}_{i,\nu_{\alpha}(f_{i})+1}\cap\mathbf{A}^{2}_{m}\subset\mathcal{H}^{\gamma}_{m}$
However, $x^{(\alpha_{1})}\neq 0$ in $U$ and $x^{(\alpha_{1})}=0$ in $\mathcal{H}^{\gamma}_{m}$, which is a contradiction. So the inclusion if it occurs is always strict. Now, let’s show the equivalence in the theorem: Suppose that $m=\nu_{\alpha}(f)$ and $\alpha=\beta$. Then,
$\overline{F^{\alpha}_{i,\nu_{\alpha}(f_{i})}\cap\mathbf{A}^{2}_{m}}=\{x^{(0)}=...=x^{(\alpha_{1}-1)}=y^{(0)}=...=y^{(\alpha_{2}-1)}=f_{i}^{[\nu_{\alpha}(f_{i})]}=0\}$
with $f^{[\nu_{\alpha}(f_{i})]}_{i}=[{y^{(\alpha_{2})}}]^{p}-c_{i}\cdot[{x^{(\alpha_{1})}}]^{q}\mod\alpha$ where $T(f_{i})=\mathbf{R}_{\geq 0}(p,q)$. Then, we have the inclusion: $\mathcal{H}^{\gamma}_{m}\subseteq\overline{F^{\alpha}_{i,\nu_{\alpha}(f_{i})+m-\nu_{\alpha}(f)}\cap\mathbf{A}^{2}_{m}}$, with $m=\nu_{\alpha}(f)$ and $\gamma=\alpha+(1,1)$.
Now given that we have the inclusion above with $m\geq\nu_{\alpha}(f)$ together with the condition $\alpha=\beta$. We have:
$\mathcal{H}^{\gamma}_{m}\neq\overline{F^{\alpha}_{i,\nu_{\alpha}(f_{i})+m-\nu_{\alpha}(f)}\cap\mathbf{A}^{2}_{m}}$
Let’s compare the codimensions of these irreducible sets:
$|\gamma|=|\alpha|+2>|\alpha|+1+m-\nu_{\alpha}(f)$
Then, $1>m-\nu_{\alpha}(f)\geq 0$. This shows $m=\nu_{\alpha}(f)$.
Now, assume on the other hand $\alpha\neq\beta$ with $\mathcal{H}^{\gamma}_{m}\subseteq\overline{F^{\alpha}_{i,\nu_{\alpha}(f_{i})+m-\nu_{\alpha}(f)}\cap\mathbf{A}^{2}_{m}}$. So, $\mathcal{H}^{\gamma}_{m}\subseteq\mathcal{H}^{\alpha}_{m}$ and $\gamma\geq\alpha$ component-wise. By comparing codimensions,
$|\gamma|\geq 1+|\alpha|+m-\nu_{\alpha}(f)$
Write $\beta=b(p,q)$ and $\alpha=a(p,q)$ then $\gamma=(bp+1,bq+1)$. Using $\gamma\geq\alpha$ we find $bp+1\geq ap$ and $bq+1\geq aq$. Since $\alpha\neq\beta$ then $a\neq b$. If $a>b$ then
$bq+1\geq aq>bq$
are integers. Then $aq=bq+1$ then (ab)q=1(a-b)q=1. So, q=ab=1q=a-b=1. Using $bp+1\geq ap>bp$ we get p=q=ab=1p=q=a-b=1. So, $\beta=(b,b)$ and $\alpha=(b+1,b+1)$. Thus, $\gamma=\beta+(1,1)=\alpha$. Now, recall $\nu_{\gamma}(f)>m\geq\nu_{\alpha}(f)$ this a contradiction with $\gamma=\alpha$. So, if $\alpha\neq\beta$ then
$\mathcal{H}^{\gamma}_{m}\not\subset\overline{F^{\alpha}_{i,\nu_{\alpha}(f_{i})+m-\nu_{\alpha}(f)}\cap\mathbf{A}^{2}_{m}}$
Now suppose $\alpha\neq\beta$ with $m=\nu_{\alpha}(f)$. If we had
$\mathcal{H}^{\gamma}_{m}\subseteq\overline{F^{\alpha}_{i,\nu_{\alpha}(f_{i})}\cap\mathbf{A}^{2}_{m}}\subseteq\mathcal{H}^{\alpha}_{m}$
Then by comparing codimension, $|\gamma|\geq 1+|\alpha|$. Also we have component-wise $\gamma\geq\alpha$. Write again $\gamma=(bp+1,bq+1)$ and $\alpha=(ap,aq)$. Thus,
$b(p+q)+2\geq 1+a(p+q)$
and $b(p+q)+1\geq a(p+q)$. Since $a\neq b$ and if $a>b$ then
$b(p+q)+1\geq a(p+q)>b(p+q)$
Since we are working with integers then (ab)(p+q)=1(a-b)(p+q)=1. Thus,
ab=p=q=1a-b=p=q=1
As before, we get $\gamma=\alpha$ and a contradiction with $\nu_{\gamma}(f)>m\geq\nu_{\alpha}(f)$. ∎
Now, let’s make sure (in the next theorem) that distinct infinite components (of the jet schemes of a Newton Plane Curves) come from distinct branches of this curve.
In other terms, if one such irreducible component is inside another infinite component then these components are in fact the same.
Before that, recall the formula of the codimension of $G^{\alpha}_{k}\cap\mathbf{A}^{2}_{m}$ where $G$ is defined by an irreducible Newton non-degenerate $g$ that divides $f$
$codim(G^{\alpha}_{k}\cap\mathbf{A}^{2}_{m};\mathbf{A}^{2}_{m})=1+|\alpha|+m-\nu_{\alpha}(f)$
where we set $k=\nu_{\alpha}(g)+m-\nu_{\alpha}(f)$.
Theorem 5.
Let $f\in\mathcal{O}_{k^{2}}$ define a Newton plane curve. Let $f_{i}$ and $f_{j}$ define branches of $f$ such that
$T(f_{l})=\mathbf{R}_{\geq 0}\alpha_{l}:\alpha_{l}\in\mathbf{N}^{2}$
and $\nu_{\alpha_{l}}(f)\leq m$ for both $l\in\{i,j\}$. Set $k_{l}=\nu_{\alpha_{l}}(f_{l})+m-\nu_{\alpha_{l}}(f)$. If
$F^{\alpha_{i}}_{i,k_{i}}\cap\mathbf{A}^{2}_{m}\subseteq\overline{F^{\alpha_{j}}_{j,k_{j}}\cap\mathbf{A}^{2}_{m}}$
then $\alpha_{i}=\alpha_{j}$ and
$\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha_{i}}_{i,k_{i}}}=\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha_{j}}_{j,k_{j}}}$
In particular, $f_{i}=f_{j}$.
Proof.
The inclusion: $F^{\alpha_{i}}_{i,k_{i}}\cap\mathbf{A}^{2}_{m}\subseteq\overline{F^{\alpha_{j}}_{j,k_{j}}\cap\mathbf{A}^{2}_{m}}$ tells us that
$1+m-\nu_{\alpha_{i}}(f)+|\alpha_{i}|\geq 1+m-\nu_{\alpha_{j}}(f)+|\alpha_{j}|$
by comparing codimensions. Then, we get the inequality by simplifying:
$|\alpha_{i}-\alpha_{j}|=|\alpha_{i}|-|\alpha_{j}|\geq\nu_{\alpha_{i}}(f)-\nu_{\alpha_{j}}(f)$ (3)
As $U:=\mathbf{A}^{2}_{m}\cap F^{\alpha_{i}}_{i,k_{i}}\subseteq\mathcal{H}^{\alpha_{j}}_{m}$ and $x^{(\alpha_{i,1})}\neq 0$ in $U$ then $\alpha_{i,1}\geq\alpha_{j,1}$ and similarly $\alpha_{i,2}\geq\alpha_{j,2}$ using $y^{(\alpha_{i,2})}\neq 0$ in $U$. Choose $\eta\in supp(f)$ such that $\eta\cdot\alpha_{i}=\nu_{\alpha_{i}}(f)$ and set $\delta=\alpha_{i}-\alpha_{j}\in\mathbf{N}^{2}_{0}$. Since $\eta\in supp(f)$ then $\eta\cdot\alpha_{j}\geq\nu_{\alpha_{j}}(f)$.
Then,
$|\delta|=|\alpha_{i}-\alpha_{j}|\geq\nu_{\alpha_{i}}(f)-\nu_{\alpha_{j}}(f)\geq\eta\cdot\alpha_{i}-\eta\cdot\alpha_{j}=\eta\cdot\delta$
So, $|\delta|\geq\nu_{\delta}(f)$. Choose $(a,b)\in supp(f)$ such that $\nu_{\delta}(f)=a\delta_{1}+b\delta_{2}\leq\delta_{1}+\delta_{2}$. This implies that $a=0$ or $b=0$ but not both, since $f(0)=0$. Since $a$ and $b$ play symmetric roles then suppose $a=0$ with $b\neq 0$. So, $b\delta_{2}\leq\delta_{1}+\delta_{2}$. Consider two cases:
  • If $\delta_{2}>\delta_{1}$ then $2\delta_{2}>\delta_{1}+\delta_{2}\geq b\delta_{2}$ thus $2>b$ and $b=1$. Then $(0,1)\in supp(f)$. Since we can take $f$ to be Weierstrass with respect to $y$, then, $f=y+a(x)$. $\frac{\partial f}{\partial y}=1$, then $(f=0)$ is not singular at the origin. Hence the contradiction.
  • If $0<\delta_{2}\leq\delta_{1}$, let’s get back to $|\delta|\geq\eta\cdot\delta$. Thus, $\delta_{1}(1-\eta_{1})+\delta_{2}(1-\eta_{2})\geq 0$.
    $(2-\eta_{1}-\eta_{2})\delta_{1}\geq 0$
    Then $\eta_{1}+\eta_{2}=2$ then $\eta_{1}=\eta_{2}=1$ since both are $\geq 1$. Since $\nu_{\alpha_{i}}(f)=\eta\cdot\alpha_{i}$, this shows $\nu_{\alpha_{i}}(f)=|\alpha_{i}|$. Then, $\alpha_{i}\cdot[\eta-(1,1)]=0$. Then, $\alpha_{i}\perp\eta-(1,1)$. Moreover, if $\eta_{1}=0$ or $\eta_{2}=0$ then $in_{\alpha_{i}}(f)$ is a monomial, which is absurd with the fact that $\alpha_{i}\in T(f)$. So, $\eta\geq(1,1)$. Now, as $\alpha_{i}$ and $\eta-(1,1)$ are both in the same quadrant $\mathbf{R}^{2}_{\geq 0}$ they can’t be orthogonal. Hence the contradiction.
  • Finally if $\delta_{1}\geq\delta_{2}=0$. Using $|\delta|\geq\eta\cdot\delta$, then again:
    $(2-\eta_{1}-\eta_{2})\delta_{1}\geq 0$
    If $\delta_{1}>0$ then $\eta_{1}=\eta_{2}=1$ since both are $\geq 1$. We find in the same way as the previous point a contradiction. So, $\delta_{1}=0$ and thus, $\delta_{1}=\delta_{2}=0$.
We found $\delta=(0,0)$. So, $\alpha_{i}=\alpha_{j}$. Now computing codimensions of the inclusion:
$\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha_{i}}_{i,k_{i}}}\subseteq\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha_{j}}_{j,k_{j}}}$
we find that they have the same codimension (since $\alpha_{i}=\alpha_{j}$). Since the parts of the inclusion are irreducible with same codimension then they are equal. This completes the proof. ∎
Below, in the next lemmas and theorem, we show that $\mathcal{H}^{\gamma}_{m}$ and $\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha}_{i,k}}$ are not comparable by inclusion, defining then irreducible components of $\mathcal{C}^{0}_{m}$.
Lemma 4.
Let $f\in\mathcal{O}_{k^{2}}$ define a Newton plane curve. Let $f_{i}$ define a branche of $f$ such that
$T(f_{i})=\mathbf{R}_{\geq 0}\alpha_{i}:\alpha_{i}\in\mathbf{N}^{2}$
and $\nu_{\alpha_{i}}(f)\leq m$. Set $k_{i}=\nu_{\alpha_{i}}(f_{i})+m-\nu_{\alpha_{i}}(f)$ and let $\gamma\in\mathbf{N}^{2}$ such that $\nu_{\gamma}(f)>m$ with $\gamma$ minimal for the product order. Then, we cannot have
$\mathcal{H}^{\gamma}_{m}\supseteq\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha_{i}}_{i,k_{i}}}$
Proof.
As $\mathbf{A}^{2}_{m}\cap F^{\alpha_{i}}_{i,k_{i}}\subseteq\mathcal{H}^{\gamma}_{m}$ then $\alpha_{i}\geq\gamma$ for the product order. Then,
$m<\nu_{\gamma}(f)\leq\nu_{\alpha_{i}}(f)\leq m$
Hence the contradiction. This concludes the proof. ∎
Lemma 5.
Let $f\in\mathcal{O}_{k^{2}}$ define a Newton plane curve. Let $f_{i}$ define a branche of $f$ such that
$T(f_{i})=\mathbf{R}_{\geq 0}\alpha_{i}:\alpha_{i}\in\mathbf{N}^{2}$
and set $\nu_{\alpha_{i}}(f)=m$, $k_{i}=\nu_{\alpha_{i}}(f_{i})$ and $\gamma=\alpha_{i}+(1,1)\in\mathbf{N}^{2}$. Then, we have
$\mathcal{H}^{\gamma}_{m}\subseteq\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha_{i}}_{i,k_{i}}}$
Moreover, if $m>\nu_{\alpha_{i}}(f)$ then $\mathcal{H}^{\gamma}_{m}\not\subset\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha_{i}}_{i,k_{i}+h}}$ where $h=m-\nu_{\alpha_{i}}(f)$.
Proof.
When, $m=\nu_{\alpha_{i}}(f)$, this comes from the fact that
$\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha_{i}}_{i,k_{i}}}=\mathcal{H}^{\alpha_{i}}_{m}\cap\{f^{(k_{i})}_{i}=0\}$
is irreducible with codimension $|\alpha_{i}|+1$, the same as $\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha_{i}}_{i,k_{i}}}$. Now, suppose $m>\nu_{\alpha_{i}}(f)$ and $\mathcal{H}^{\gamma}_{m}\subsetneq\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha_{i}}_{i,k_{i}+h}}$. Then, by comparing codimensions:
$|\alpha_{i}|+2>|\alpha_{i}|+1+h$
Thus, $1>h=m-\nu_{\alpha_{i}}(f)>0$. Hence the contradiction. ∎
Theorem 6.
Let $f\in\mathcal{O}_{k^{2}}$ define a Newton plane curve. Let $f_{i}$ define a branche of $f$ such that
$T(f_{i})=\mathbf{R}_{\geq 0}\alpha:\alpha\in\mathbf{N}^{2}$
and $\nu_{\alpha}(f)\leq m$. Set $k=\nu_{\alpha}(f_{i})+m-\nu_{\alpha}(f)$. Let $\gamma\in\mathbf{N}^{2}$ be minimal for the product order, such that $\nu_{\gamma}(f)>m$ and $\gamma\neq\alpha+(1,1)$. Then, we have
$\mathcal{H}^{\gamma}_{m}\not\subset\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha}_{i,k}}$
Proof.
Suppose (to find a contradiction) the inclusion holds. Then $\gamma\geq\alpha$ component-wise. Let’s show $|\gamma|\leq 1+|\alpha|+m-\nu_{\alpha}(f)$ by induction on $m\geq\nu_{\alpha}(f)$.
  • if $m=\nu_{\alpha}(f)$ then by the minimality of $\gamma\geq\alpha$ we have either $\gamma=\alpha+(1,0)$ or $\gamma=\alpha+(0,1)$ because if we set $n_{i}=\nu_{\alpha}(f_{i})$, $r=\nu_{\alpha}(R)$ and $f=Rf_{i}$
    $f^{(m)}=f^{(n_{i})}_{i}R^{(r)}\mod\alpha$
    with $Rad\langle R^{(r)}\rangle=\langle x^{(\alpha_{1})}y^{(\alpha_{2})}\cdot*\rangle$ where * is a product of pairwise distinct binomials. So this shows $|\gamma|=|\alpha|+1$.
  • if $m>\nu_{\alpha}(f)$ and $\nu_{\gamma}(f)>m+1$ then $\gamma$ is still minimal at level $m+1$ and
    $|\gamma|\leq 1+|\alpha|+m+1-\nu_{\alpha}(f)$
  • if $m>\nu_{\alpha}(f)$ and $\nu_{\gamma}(f)=m+1$ then $\gamma$ (which is minimal) either bifurcates to $\gamma+(1,0)$ or $\gamma+(0,1)$ or gives rise to a component when $\gamma\in T(f)$
    $\overline{\mathbf{A}^{2}_{m}\cap F^{\gamma}_{j,\nu_{\gamma}(f_{j})+1}}$
    whose codimension is $|\gamma|+1$. If we are in the first or second case, since $\nu_{\gamma}(f)<\nu_{\gamma+\epsilon}(f)$ then at level $m+2$ we have:
    $|\gamma|+1\leq 1+|\alpha|+m+2-\nu_{\alpha}(f)$
    with $\gamma+\epsilon$ minimal. The last case when $\gamma\in T(f)$ tell us that
    $\mathbf{A}^{2}_{m}\cap F^{\gamma}_{j,\nu_{\gamma}(f_{j})+1}\subset\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha}_{i,k}}$
    which contradicts the previous theorem.
So, we have shown $|\gamma|\leq 1+|\alpha|+m-\nu_{\alpha}(f)$. Together with the inclusion which we supposed we find a clear contradiction (unless $\gamma=\alpha+(1,1)$). The case $\gamma=\alpha+(1,1)$ is excluded by hypothesis. ∎
7. Main Theorems
In this section, we describe the irreducible components of the jet schemes over the origin of a Newton non-degenerate plane curve singularity. First, let’s present some definitions and notations.
Let $f\in\mathcal{O}_{k^{2}}$ be Newton non-degenerate. A tropical decomposition of $f$ is $f=f_{1}...f_{t}$ where $T(f_{i})$ is a semi-line in $\mathbf{R}_{\geq 0}^{2}$ passing through the origin and $T(f_{i})\cap T(f_{j})=\{(0,0)\}$ for $1\leq i<j\leq t$.
Denote now by $\sigma=Conv[T(f)]$. Introduce the notations
$A_{m}(i)=\{\alpha\in T(f_{i})\cap\mathbf{N}^{2}:\nu_{\alpha}(f)\leq m\}$
$B^{*}_{m}\subseteq B_{m}=\min_{\leq\times\leq}\{\gamma\in\mathbf{N}^{2}:\nu_{\gamma}(f)>m\}$
where $B^{*}_{m}$ is obtained by removing from $B_{m}$ at $m=\nu_{\alpha}(f)$ every $\alpha+(1,1)$ : $\alpha\in B_{m-1}\cap\partial\sigma$. Theorem 7 gives us the decomposition of $\mathcal{C}^{0}_{m}$ into irreducible components.
Theorem 7.
Let $f=f_{1}...f_{t}\in\mathcal{O}_{k^{2}}$ be a tropical decomposition of a Newton non-degenerate polynomial. Decompose $f_{i}=\prod_{j=1}^{r_{i}}f_{(i,j)}$ into irreducible factors in $\mathcal{O}_{k^{2}}$. Then, we have the decomposition into irreducible components
$\mathcal{C}^{0}_{m}=\bigcup_{\gamma\in B^{*}_{m}}\mathcal{H}^{\gamma}_{m}\cup\bigcup_{i=1}^{t}\bigcup_{\alpha\in A_{m}(i)}\bigcup_{j=1}^{r_{i}}\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha}_{(i,j),k_{ij}}}$
where $k_{ij}=\nu_{\alpha}(f_{ij})+m-\nu_{\alpha}(f)$.
Proof.
Recall for $v\in\{x,y\}$: $Cont^{>0}_{\mathcal{C}}(v)_{m}=\bigsqcup_{i=1}^{m}Cont^{i}_{\mathcal{C}}(v)_{m}$ where we set
$Cont^{i}_{\mathcal{C}}(v)_{m}=\{v^{(i)}\neq 0\}\cap\{v^{(0)}=...=v^{(i-1)}=0\}\cap\{f^{(0)}=...=f^{(m)}=0\}$
Set also, $Cont^{(i,j)}_{\mathcal{C}}(x,y)_{m}=Cont^{i}_{\mathcal{C}}(x)_{m}\cap Cont^{j}_{\mathcal{C}}(y)_{m}$. Start with the observation,
$\mathcal{C}^{0}_{m}=\mathcal{H}^{(1,1)}_{m}\cap F_{m}=Cont^{>0}_{\mathcal{C}}(x)_{m}\cap Cont_{\mathcal{C}}^{>0}(y)_{m}=\bigsqcup_{1\leq i,j\leq m}Cont^{(i,j)}_{\mathcal{C}}(x,y)_{m}$
Then, if we set $\alpha=(i,j)$ and notice that $Cont^{(i,j)}_{\mathcal{C}}(x,y)_{m}=F^{(i,j)}_{m}\cap\mathbf{A}^{2}_{m}$, thus,
$\mathcal{C}^{0}_{m}=\bigsqcup_{\alpha\in\Omega}(F^{\alpha}_{m}\cap\mathbf{A}^{2}_{m})=\left(\bigsqcup_{\alpha\in\Omega^{>}}(F^{\alpha}_{m}\cap\mathbf{A}^{2}_{m})\right)\sqcup\left(\bigsqcup_{\alpha\in\Omega^{\leq}}(F^{\alpha}_{m}\cap\mathbf{A}^{2}_{m})\right)$
where $\Omega=[1,m]^{2}\cap\mathbf{N}^{2}$, $\Omega^{>}=\{\alpha\in\Omega:\nu_{\alpha}(f)>m\}$ and $\Omega^{\leq}=\Omega\backslash\Omega^{>}$. Set
$A_{m}=\bigcup_{i=1}^{t}A_{m}(i)$
If $\alpha\notin A_{m}$ then $f^{[\nu_{\alpha}(f)]}=c\cdot{x^{(\alpha_{1})}}^{k_{1}}{y^{(\alpha_{2})}}^{k_{2}}\mod\alpha$ for some constant $c$. As $x^{(\alpha_{1})}$ and $y^{(\alpha_{2})}$ are $\neq 0$ in $F^{\alpha}_{m}$ then $F^{\alpha}_{m}\cap\mathbf{A}^{2}_{m}=\varnothing$ if $\alpha\notin A_{m}$. Now, we can write:
$\mathcal{C}^{0}_{m}=\bigsqcup_{\alpha\in\Omega^{>}}(F^{\alpha}_{m}\cap\mathbf{A}^{2}_{m})\sqcup\bigsqcup_{i=1}^{t}\bigsqcup_{\alpha\in A_{m}(i)}(F^{\alpha}_{m}\cap\mathbf{A}^{2}_{m})$
Now if $\alpha\in A_{m}(i)$ then $F^{\alpha}_{m}=F^{\alpha}_{i,k}$ where $k=\nu_{\alpha}(f_{i})+m-\nu_{\alpha}(f)$,
$F^{\alpha}_{i,k}=\bigsqcup_{j=1}^{r_{i}}F^{\alpha}_{(i,j),k_{ij}}$
because of lemma 2. Now, as $\mathcal{H}^{\alpha}_{m}=\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha}_{m}}$ for $\nu_{\alpha}(f)>m$ then we obtain after ordering along $\leq\times\leq$ and using theorem 4,
$\mathcal{C}^{0}_{m}=\bigcup_{\gamma\in B^{*}_{m}}\mathcal{H}^{\gamma}_{m}\cup\bigcup_{i=1}^{t}\bigcup_{\alpha\in A_{m}(i)}\bigcup_{j=1}^{r_{i}}\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha}_{(i,j),k_{ij}}}$
The irreducibility of these components follow from 5, 6 and 4. ∎
We have just proven that $Jet(f)$ has two kinds of vertices: vertices of the form $\mathcal{H}^{\gamma}_{m}$ and others that do not vanish and continue to $\infty$. The below theorem characterizes them.
Theorem 8.
Let $f=f_{1}...f_{t}\in\mathcal{O}_{k^{2}}$ be Newton non-degenerate ; $T(f_{i})\simeq\mathbf{R}_{\geq 0}$,
$T(f_{i})\cap T(f_{j})=\{(0,0)\}:i<j$
Decompose $f_{i}=\prod_{j=1}^{r_{i}}f_{ij}$ into irreducible factors. Let $V_{m}\in irr(\mathcal{C}^{0}_{m})$ with weight $(d,e)$. Then, $V_{m}$ is of the form $\mathcal{H}^{\gamma}_{m}$ for some $\gamma\in B^{*}_{m}$ if and only if $e=d$. Otherwise, $e>d$.
Proof.
In this proof, set $c=codim(V_{m};\mathbf{A}^{2}_{m})$. If $V_{m}=\mathcal{H}^{\gamma}_{m}$ then $c=|\gamma|$. Now, let’s show $e+c=2(m+1)$.
$\mathcal{O}_{V_{m}}=\mathcal{O}_{\mathbf{A}^{2}_{m}}/\langle x^{(0)},...,x^{(\gamma_{1}-1)},y^{(0)},...,y^{(\gamma_{2}-1)}\rangle$
$e\in\mathbf{N}$ is the least integer such that we have a surjective map of rings
$\mathcal{O}_{k^{e}}\rightarrow\mathcal{O}_{V_{m}}$
Thus, $e=dim(\mathbf{A}^{2}_{m})-|\gamma|=2(m+1)-|\gamma|$. Thus, $e+c=2(m+1)$ ie $e=d$. Now, if $V_{m}$ isn’t of the form $\mathcal{H}^{\gamma}_{m}$ then there exists $\alpha\in T(f_{ij})$ for some $i,j$ such that
$V_{m}=\overline{\mathbf{A}^{2}_{m}\cap F^{\alpha}_{ij,\nu_{\alpha}(f_{ij})+m-\nu_{\alpha}(f)}}$
with $m\geq\nu_{\alpha}(f)$. Also, $c=|\alpha|+1+m-\nu_{\alpha}(f)$, there are surjective ring maps:
$\mathcal{O}_{k^{e}}\rightarrow\mathcal{O}_{V_{m}}\rightarrow\mathcal{O}_{\mathcal{H}^{\alpha}_{m}}$
Thus, $e\geq 2(m+1)-|\alpha|$. Then, we obtain:
$e+c\geq 2(m+1)-|\alpha|+c=2(m+1)+1+m-\nu_{\alpha}(f)$
Thus, $e+c>2(m+1)$. Thus, $e>d$. This completes the proof.
Now, we show that $Jet(f)$ determines the embedded topological type of the singularity of the Newton non-degenerate plane curve defined by $f$.
Theorem 9.
Let $f,g\in\mathcal{O}_{k^{2}}$ be Newton non-degenerate. If the jet-components graph $Jet(f)=Jet(g)$ then $f$ and $g$ have the same embedded topological type.
Proof.
According to theorem 8, $J:=Jet(f)$ determines $B^{*}:=\bigcup_{m\in\mathbf{N}_{0}}B^{*}_{m}$. In other words, $B^{*}_{m}:=B^{*}_{m}(f)=B^{*}_{m}(g)$ for all $m$. Denote by $R^{*}$ the induced subgraph of $J$ obtained by removing the vertices of the form $\mathcal{H}^{\gamma}_{m}$ : $\gamma\in B^{*}_{m}$, $\nu_{\gamma}(f)>m\in\mathbf{N}_{0}$ and $\mathcal{H}^{\gamma}_{m}$ could not be joined with a vertex outside of $B^{*}$. Define $A=\{\alpha:\exists m:\mathcal{H}^{\alpha}_{m}\in R^{*}\}$. Note that:
$\alpha\in A\Leftrightarrow\mathcal{H}^{\alpha}_{\nu_{\alpha}(f)-1}\in R^{*},\alpha\in T(f)\cap\mathbf{N}^{2}$
Set $T=\bigcup_{\alpha\in A}\mathbf{R}_{\geq 0}\cdot\alpha$ which is then the tropical variety of both $f$ and $g$. Now, let’s denote $f=f_{1}...f_{r}$ and $g=g_{1}...g_{s}$ as product of branches. Let $\alpha_{i}\in A$. Denote by $T_{\alpha_{i}}=\mathbf{R}_{\geq 0}\cdot\alpha_{i}$ the tropical variety of the branches $f_{1},...,f_{r_{i}},g_{1},...,g_{s_{i}}$ with $r_{i}\leq r$ and $s_{i}\leq s$. Since $f$ and $g$ are Newton non-degenerate with the same graph $J$ then the there are exactly $r_{i}=s_{i}$ vertices above $\mathcal{H}^{\alpha_{i}}_{\nu_{\alpha_{i}}(f)-1}=\mathcal{H}^{\alpha_{i}}_{\nu_{\alpha_{i}}(g)-1}$ connected to it. Thus, $r_{i}=s_{i}$ and $r=s$. Also, for $1\leq i<j\leq l<\ell\leq r$,
$(f_{i},f_{j})_{0}=\nu_{\alpha_{i}}(f_{i})=\nu_{\alpha_{i}}(f_{j})$
$(f_{i},f_{\ell})_{0}=\nu_{\alpha_{i}}(f_{\ell})$
The same for $g$. Denote by $F=f/(f_{1}...f_{l})$ and $G=g/(g_{1}...g_{l})$. So,
$l\nu_{\alpha_{i}}(f_{1})+\nu_{\alpha_{i}}(F)=\nu_{\alpha_{i}}(f)=\nu_{\alpha_{i}}(g)=l\nu_{\alpha_{i}}(g_{1})+\nu_{\alpha_{i}}(G)$
$\nu_{\alpha_{i}}(f_{1})=lcm(\alpha_{i,1},\alpha_{i,2})=\nu_{\alpha_{i}}(g_{1})$. Since $\nu_{\alpha_{i}}(f)=\nu_{\alpha_{i}}(g)$ then $\nu_{\alpha_{i}}(F)=\nu_{\alpha_{i}}(G)$. This holds for all $\alpha_{i}\in A$. So, we have shown that $f$ and $g$ have the same number of branches, the same intersection multiplicities (of their branches) and the same tropical varieties for branches (since $r_{i}=s_{i}$). Now using theorem 3, $f$ and $g$ have the same embedded topological type. ∎
As we did in Figures 1,3 and 4, we can associate to $J=Jet(f)$ a special staircase walk $J_{SC}(f)$ in $\mathbf{R}^{2}_{\geq 0}$. Retake the same notations of theorem 9. Denote by $T_{i}=T(f_{i})$ where we suppose $T_{j}$ above $T_{i}$ for $i<j$. Let’s define a dictionnary for this walk: every drawn $\alpha\in\mathbf{N}^{2}$, represents the irreducible components of the form $\mathcal{H}^{\alpha}_{m}$ of $\mathcal{C}^{0}_{m}$ : $m<\nu_{\alpha}(f)$. The red arrows correspond to irreducible components of $\mathcal{C}^{0}_{m}$ that don’t vanish and continue till $\infty$.
  1. (1)
    If $t=1$, then the staircase walk is identical to that of Figure 3 except that at each $\alpha\in T(f)$ we add $r$ red arrows above $\alpha$.
    Here is a more detailed description of it: let $\lambda\in L_{(p,q)}\cap\mathbf{N}^{2}_{0}$ and recall the staircase notation $d_{1},...,d_{p}]]$ associated to $(p,q)$ from the theorem 1:
    • Start at $(1,1)$ ie $\lambda=(0,0)$ here;
    • Move $d_{1}$ vertical unit steps followed by one horizontal step;
    • Move $d_{2}$ vertical unit steps followed by one horizontal step;
    • Move $d_{p-1}$ vertical unit steps followed by one horizontal step;
    • Finally, move $d_{p}-1$ vertical steps; we reach $(p,q)$;
    • Add this time $r$ red arrow above $(p,q)$;
    • Take $\lambda=(p,q)$ and start again;
    Repeat the above steps whenever we reach $\lambda\in L_{(p,q)}\cap\mathbf{N}_{0}^{2}$ till reaching $\lambda+(p,q)$ above which we add $r$ red arrows.
    Note that each time we move from a point $\alpha$ to $\alpha+(0,1)$ or $\alpha+(1,0)$, we delete $\alpha$ and we replace it by the new point. The red arrows are never deleted.
  2. (2)
    If $t=2$, then set $\sigma=T_{1}+T_{2}$; the staircase walk obeys to the rules:
    • Start always at $(1,1)$;
    • If $\alpha\in\overset{\circ}{\sigma}$, remove $\alpha$ to replace it with $\alpha+(0,1)$ and $\alpha+(1,0)$;
    • If $\alpha\in T_{1}$, add $r_{1}$ red arrows above $\alpha$ and add $\alpha+(1,0)$;
    • If $\alpha$ is above $T_{2}$ replace $\alpha$ by $\alpha+(1,0)$;
    • If $\alpha\in T_{2}$, add $r_{2}$ red arrows above $\alpha$ and add $\alpha+(0,1)$;
    • If $\alpha$ is below $T_{1}$ replace $\alpha$ by $\alpha+(0,1)$.
  3. (3)
    If $t>2$, set $T=T(f)$, $\sigma=T_{1}+T_{t}$; the staircase walk obeys to the rules:
    • Start always at $(1,1)$;
    • If $\T\alpha\in\sigma\backslash$ T, remove $\alpha$ to replace it with $\alpha+(0,1)$ and $\alpha+(1,0)$;
    • If $\alpha\in T_{i}$ : $2\leq i\leq t-1$ add $r_{i}$ red arrows above $\alpha$ and add $\alpha+(1,0)$ with $\alpha+(0,1)$;
    • If $\alpha\in T_{1}$, add $r_{1}$ red arrows above $\alpha$ and add $\alpha+(1,0);$
    • If $\alpha$ is above $T_{t}$ replace $\alpha$ by $\alpha+(1,0)$;
    • If $\alpha\in T_{t}$, add $r_{t}$ red arrows above $\alpha$ and add $\alpha+(0,1)$;
    • If $\alpha$ is below $T_{1}$ replace $\alpha$ by $\alpha+(0,1)$.
Irreducible components Interpretation Jet-components graph
$\mathcal{H}^{\alpha}_{m}:m<\nu_{\alpha}(f)$ Big, Vanishing $\alpha,\nu_{\alpha}(f)-1]$
$\mathbf{A}^{2}_{m}\cap F^{\alpha}_{ij,k}:\alpha\in T(f_{i})$ Nonvanishing $(\alpha,ij,\infty)$
Table 3. Dictionary of notations for $f$ Newton non-degenerate.
Theorem 10.
For $\mathcal{C}=\{f=0\}\subset k^{2}$ a Newton plane curve, we can extract the irreducible components of $\mathcal{C}^{0}_{m}$ for all $m\in\mathbf{N}$ from $J_{SC}(f)$.
Proof.
Consider first an edge and denote by $(\alpha,w)$ and $(\alpha^{\prime},w^{\prime})$ with $w<w^{\prime}<\infty$ the weighted vertices of this edge. Note that $\alpha^{\prime}=\alpha+(1,0)$ or $\alpha^{\prime}=\alpha+(0,1)$ with $w=\nu_{\alpha}(f)-1$ and $w^{\prime}=\nu_{\alpha^{\prime}}(f)-1$. Add on the edge $\alpha\alpha^{\prime}$, $w^{\prime}-w-1$ vertices representing the big components defined by the same $\alpha$. Now consider a nonvanishing component represented by $(\alpha,ij,\infty)$. Add on the arrow above $\alpha$ a countably infinite amount of vertices. These represent the nonvanishing components for $m\geq\nu_{\alpha}(f)$ and $\alpha\in T(f)$. Finally, scale each weight $w$ by its level $m=w+1$. The obtained graph is the jet components graph. ∎
8. Generating Series
In this section, we introduce the generating series associated to $f\in K[[x,y]]$ a Newton non-degenerate power series. Denote by $\mathcal{C}=\{f=0\}$ its associated plane curve singularity. This series will be denoted $G:=G(f)$.
$G=\sum_{K,m}u^{\text{codim}(K;\mathbf{A}^{2}_{\infty})}\cdot v^{m}\in\mathbf{Z}[[u,v]]$
where $m\geq 1$ and $K$ is an irreducible component in $\text{Cont}^{m}(\mathcal{C})^{0}\neq\varnothing$. We call a subsummation of a sum $S$ to be a sum whose terms form a subset of the terms of $S$.
This section is dedicated to showing that $G=G(f)$ is rational in $u,v$, this means that $G\in\mathbf{Z}(u,v)$. We shall always write $G=H+R$ where
  • $H$ is the subsummation of $G$ with $K$ being of the form $\mathcal{H}^{\beta}_{\infty}$;
  • $R$ is the rest of the sum since $G=H+R$.
We used the letter $H$ because the $K$s appearing in $H$ are all irreducible components having hyperplane coordinates. R=GHR=G-H is just the rest of the sum.
Let’s consider two different examples at first.
Example 7.
Let’s start with the example of $f=(y^{2}-x^{3})(y^{2}-2x^{3})$. Its tropical variety $T(f)=\mathbf{R}_{\geq 0}(2,3)$ is a semi-line. Set $G=G(f)$. Introduce
$R:=2\sum_{\alpha,m}u^{1+|\alpha|+m-\nu_{\alpha}(f)}v^{m}$
where $m\geq\nu_{\alpha}(f)$ and $\alpha\in T(f)$. Then, since $\alpha=s\cdot(2,3)$ for $s\in\mathbf{N}$,
$R=2u\sum_{s=1}^{\infty}\sum_{m=12s}^{\infty}u^{5s+m-12s}v^{m}=\frac{2u}{1-uv}\sum_{s=1}^{\infty}u^{5s}v^{12s}=\frac{2u}{1-uv}\frac{u^{5}v^{12}}{1-u^{5}v^{12}}$
Now, set $S_{1}=(0,4)$ and $S_{2}=(6,0)$. Introduce $S^{*}_{i}$ the corresponding cone to $S_{i}$ in the dual Newton fan. Set
$H:=\sum_{s=1}^{\infty}u^{5s}v^{12s}+\sum_{\beta_{1}}u^{|\beta_{1}|}v^{\beta_{1}\cdot S_{1}}+\sum_{\beta_{2}}u^{|\beta_{2}|}v^{\beta_{2}\cdot S_{2}}$
where $\beta_{i}\in B^{*}_{m}\cap\overset{\circ}{S^{*}_{i}},m\geq 1$. Let’s describe where $\beta_{i}$ lies.
$\beta_{1}-\lambda\in\{(1,1);(2,2)\}$
for $\lambda\in T(f)\cap\mathbf{N}_{0}^{2}$. So we get, $\beta_{2}-\lambda\in\{(1,2)\}$ for the same $\lambda=(2s,3s)$.
$H=\frac{u^{5}v^{12}}{1-u^{5}v^{12}}+\sum_{s\geq 0}u^{5s+2}v^{12s+4}+\sum_{s\geq 0}u^{5s+4}v^{12s+8}+\sum_{s\geq 0}u^{5s+3}v^{12s+6}$
Then, $H\in\mathbf{Z}(u,v)$. So, $G=H+R\in\mathbf{Z}(u,v)$.
Example 8.
Consider $f=f_{1}f_{2}=(y^{2}-x^{3})(y^{3}-x^{5})$. Its tropical variety is
$T(f)=\mathbf{R}_{\geq 0}(2,3)\cup\mathbf{R}_{\geq 0}(3,5)$
Let’s compute
$R=\underbrace{\sum_{\alpha_{1},m_{1}}u^{1+|\alpha_{1}|+m_{1}-\nu_{\alpha_{1}}(f)}v^{m_{1}}}_{=R_{1}}+\underbrace{\sum_{\alpha_{2},m_{2}}u^{1+|\alpha_{2}|+m_{2}-\nu_{\alpha_{2}}(f)}v^{m_{1}}}_{=R_{2}}$
where $\alpha_{i}\in T(f_{i})$, $m_{i}\geq\nu_{\alpha_{i}}(f)$.
$R_{1}=u\sum_{s=1}^{\infty}\sum_{m=0}^{\infty}u^{5s+m}v^{m+15s}=\frac{u}{1-uv}\frac{u^{5}v^{15}}{1-u^{5}v^{15}}$
$R_{2}=\frac{u}{1-uv}\frac{u^{8}v^{24}}{1-u^{8}v^{24}}$
Denote now, $S_{1}=(0,5)$, $S_{2}=(3,3)$ and $S_{3}=(8,0)$ the vertices of the Newton polygon from left to right. Introduce again $S_{i}^{*}$ the corresponding cone to $S_{i}$ in the dual Newton fan. Let’s compute
$H:=\sum_{m,\beta}u^{|\beta|}v^{m}=\underbrace{\sum_{r,s\geq 1}u^{5r+8s}v^{[r(2,3)+s(3,5)]\cdot(3,3)}}_{=H_{\overset{\circ}{\sigma_{2}}}}+\tilde{H}=\tilde{H}+\sum_{r,s}u^{5r+8s}v^{15r+24s}$
where $\beta\in B^{*}_{m},m\geq 1$ and $\sigma_{2}=\mathbf{R}_{\geq 0}(2,3)+\mathbf{R}_{\geq 0}(3,5)=S^{*}_{2}$. We obtain by computing geometric series
$D:=H-\tilde{H}=\frac{u^{5}v^{15}}{1-u^{5}v^{15}}\frac{u^{8}v^{24}}{1-u^{8}v^{24}}\in\mathbf{Z}(u,v)$
Now, let’s compute
$\tilde{H}=\sum_{s\geq 1}u^{5s}v^{s(2,3)\cdot(0,5)}+\sum_{s\geq 1}u^{8s}v^{s(3,5)\cdot(8,0)}+H^{*}$
where $H^{*}$ is the subsummation of $H$ whose $\beta$ lie in the first and last cone determined by orientation of the tropical decomposition.
Then, $\tilde{H}-H^{*}=\frac{u^{5}v^{15}}{1-u^{5}v^{15}}+\frac{u^{8}v^{24}}{1-u^{8}v^{24}}\in\mathbf{Z}(u,v)$. So, we have
$H^{*}=\sum_{\beta_{1}}u^{|\beta_{1}|}v^{\beta_{1}\cdot S_{1}}+\sum_{\beta_{3}}u^{|\beta_{3}|}v^{\beta_{3}\cdot S_{3}}$
where $\beta_{i}\in\overset{\circ}{S_{i}^{*}}\cap B^{*}_{m},m\geq 1$, $S_{1}=(0,5)$ and $S_{3}=(8,0)$. Then,
$\beta_{1}-\lambda_{1}\in\{(1,1),(2,2)\}:\lambda_{1}=s_{1}(2,3),s_{1}\in\mathbf{N}_{0}$
and $\beta_{3}-\lambda_{3}\in\{(1,2),(2,4)\}:\lambda_{3}=s_{3}(3,5),s_{3}\in\mathbf{N}_{0}$. So we get $H^{*}$ equal to
$\sum_{s_{1}\geq 0}(u^{5s_{1}+2}v^{15s_{1}+5}+u^{5s_{1}+4}v^{15s_{1}+10})+\sum_{s_{3}\geq 0}(u^{8s_{3}+3}v^{24s_{3}+8}+u^{8s_{3}+6}v^{24s_{3}+16})$
This leads to
$H^{*}=\frac{u^{2}v^{5}+u^{4}v^{10}}{1-u^{5}v^{15}}+\frac{u^{3}v^{8}+u^{6}v^{16}}{1-u^{8}v^{24}}\in\mathbf{Z}(u,v)$
So, $H\in\mathbf{Z}(u,v)$ and $G=H+R$ as well.
In the below lemma, we write an expression for $R$ (which was introduced at the beginning of the section).
Lemma 6.
Let $f\in K[x,y]$ be Newton non-degenerate whose tropical decomposition is written as $f=f_{1}...f_{t}$. For $1\leq i\leq t$, write $f_{i}=\prod_{j=1}^{r_{i}}f_{i,j}$. Then
$R=\sum_{i,j,\alpha,m}u^{1+|\alpha|+m-\nu_{\alpha}(f)}v^{m}$
where $1\leq i\leq t$, $1\leq j\leq r_{i}$, $\alpha\in T(f_{i})$ and $m\geq\nu_{\alpha}(f)$. Also, $R\in\mathbf{Z}(u,v)$ and
$R=\sum_{i}\frac{r_{i}u}{1-uv}\cdot\frac{u^{|\alpha_{i}|}v^{\alpha_{i}\cdot S_{i}}}{1-u^{|\alpha_{i}|}v^{\alpha_{i}\cdot S_{i}}}$
where $T(f_{i})=\mathbf{R}_{\geq 0}\cdot\alpha_{i}$, $gcd(\alpha_{i,1},\alpha_{i,2})=1$ and $S_{1},...,S_{t+1}$ are the vertices of the Newton polygon of $f$ (from left to right).
Proof.
Set $R_{i}=\sum_{j,\alpha_{i},m}u^{1+|\alpha_{i}|+m-\nu_{\alpha_{i}}(f)}v^{m}$ in such a way that $R=\sum_{i}R_{i}$. Since $\nu_{\alpha}(f)=\alpha\cdot S_{i}$ (because $\alpha\in T(f_{i})$). Thus, we obtain:
$R_{i}=r_{i}\cdot\sum_{s\geq 1}\sum_{m\geq 0}u^{1+s|\alpha_{i}|+m}v^{m+s\cdot\nu_{\alpha_{i}}(f)}=\frac{r_{i}u}{1-uv}\frac{u^{|\alpha_{i}|}v^{\alpha_{i}\cdot S_{i}}}{1-u^{|\alpha_{i}|}v^{\alpha_{i}\cdot S_{i}}}$
Which completes the proof. ∎
Now let $f=f_{1}...f_{t}$ be a tropical decomposition of a Newton non-degenerate polynomial. The decomposition $f=f_{1}...f_{t}$ is oriented when for all $i<j$,
$det[T(f_{i}),T(f_{j})]>0$
ie $T(f_{j})$ is ”above” $T(f_{i})$ which we write $T(f_{j})>T(f_{i})$. Now let’s apply this new notion of oriented tropical decomposition to obtain the below observation.
Observation 2.
Let $f$ define a Newton Plane Curve singularity such that $(1,1)\notin supp(f)$. Then for every $\alpha\in T(f)$, $\nu_{\alpha}(f)>|\alpha|$ unless $\alpha\in\mathbf{R}_{\geq 0}\cdot(1,1)$ with $f$ having two branches with $T(f)=\mathbf{R}_{\geq 0}\cdot(1,1)$.
Proof.
Since we can always consider the symmetry $(x,y)\mapsto(y,x)$, let $\alpha=(m,n)$ with $m\geq n\geq 1$ and $gcd(m,n)=1$. Suppose
$\nu_{\alpha}(f_{\alpha})\leq\nu_{\alpha}(f)\leq|\alpha|$
where $f_{\alpha}$ denotes the product of branches having tropical variety $\mathbf{R}_{\geq 0}\cdot\alpha$. Let’s show $f=f_{\alpha}$. Otherwise, $t\geq 2$ where $t$ is the number of semi-lines in the tropical variety $T(f)$. Denote by $f=f_{1}...f_{t}$ an oriented tropical decomposition of $f$. We have $f_{\alpha}\in\{f_{1},...,f_{t}\}$. There exists $S=(a,b)$ a vertex in the Newton polygon of $f$ with $a,b\geq 1$ (because $t>1$) and $\nu_{\alpha}(f)=\alpha\cdot S$. Set $\eta=(1,1)-S$
$0\leq|\alpha|-\nu_{\alpha}(f)=\alpha\cdot\eta$
Note that $\alpha_{1},\alpha_{2}\geq 1$. Since $a,b\geq 1$ then $\alpha\cdot\eta=0$ but $\alpha$ is in the first positive quadrant $\mathbf{R}^{2}_{>0}$ and $\eta\in\mathbf{R}^{2}_{\leq 0}$ so they can’t be orthogonal unless $\eta=0$ ie $S=(1,1)\in supp(f)$, a contradiction. So $t=1$ and $f=f_{\alpha}$.
$mn\leq rmn\leq m+n$
where $r$ is the number of branches of $f$. According to Observation 1, $n=1$ then
$1\leq r\leq 1+\frac{1}{m}$
If $m>1$ then $r=1$ we can write in Weierstrass form
$f=y+a(x)$
Thus $f$ is smooth, a contradiction with $f$ defining a singularity. So $r=2$ and $m=1$. So, we can write in Weierstrass form after doing a change of variables
f=(yx+)(ycx+)f=(y-x+...)(y-cx+...)
Thus, $|\alpha|=\nu_{\alpha}(f)=2$ with $\alpha=(1,1)$. This concludes the proof of our observation. ∎
In the below lemma, we write an expression for $H$ (which was introduced at the beginning of the section). There are many cases to consider:
  • if $T=\cup_{i}T_{i}$ is the tropical variety of $f$, then $H_{T_{i}}$ is the subsummation of $H$ where $\beta\in T_{i}$;
  • if $\sigma_{i}=T_{i-1}+T_{i}$ then $H_{\overset{\circ}{\sigma_{i}}}$ is the subsummation of $H$ whose $\beta$ lie in the interior $\overset{\circ}{\sigma_{i}}$ of the cone.
Lemma 7.
Let $f\in K[x,y]$ be a Newton non-degenerate polynomial whose oriented tropical decomposition is denoted $f=f_{1}...f_{t}$. Then $H$ is the summation
$H=\sum_{m,\beta}u^{|\beta|}v^{m}$
where $m\geq 1$, $Cont^{m}(\mathcal{C})^{0}\neq\varnothing$ and $\beta\in B^{*}_{m}$. Also, $H\in\mathbf{Z}(u,v)$. Moreover, if $S_{1},...,S_{t+1}$ are the vertices of the Newton polygon of $f$ from left to right, we have
  • If $t=1$ and $T(f)=\mathbf{R}_{\geq 0}(1,1)$ then $H=\frac{u^{2}v^{s}}{1-u^{2}v^{s}}$ where the Newton polygon of $f$ is the segment joining $(s,0)$ and $(0,s)$,
  • If $t=1$ and $T(f)>\mathbf{R}_{\geq 0}(1,1)$ then $H=H_{T}+H_{\overset{\circ}{\sigma_{1}}}+H_{\overset{\circ}{\sigma_{2}}}$ where
    $H_{T}=\frac{u^{|\alpha|}v^{\alpha\cdot S}}{1-u^{|\alpha|}v^{\alpha\cdot S}}$
    where $T(f)=\mathbf{R}_{\geq 0}\cdot\alpha$, $gcd(\alpha_{1},\alpha_{2})=1$ and $S=(s,0)$ is a vertex of the Newton polygon of $f$,
    $H_{\overset{\circ}{\sigma_{1}}}=\frac{1}{1-u^{|\alpha|}v^{\alpha\cdot S}}\sum_{i=1}^{p}\sum_{j=D_{i-1}+1}^{D_{i}}u^{i+j}v^{js}$
    $H_{\overset{\circ}{\sigma_{2}}}=\frac{1}{1-u^{|\alpha|}v^{\alpha\cdot S}}\sum_{i=1}^{p}u^{i+D_{i}+1}v^{is}$
    where $SC(\alpha_{2}/\alpha_{1})=[[d_{1},...,d_{p}]]$ and $D_{i}=\sum_{j=1}^{i}d_{j}$, $D_{0}=0$.
  • If $t>1$, set $T_{i}=T(f_{i})=\mathbf{R}_{\geq 0}\cdot\alpha_{i}$ with $T_{1}<...<T_{t}$ and $T_{1}=\mathbf{R}_{\geq 0}(1,1)$ then
    $H=H_{T_{1}}+...+H_{T_{t}}+H_{\overset{\circ}{\sigma_{2}}}+...+H_{\overset{\circ}{\sigma_{t+1}}}$
    where $H_{T_{i}}=\frac{u^{|\alpha_{i}|}v^{\alpha_{i}\cdot S_{i}}}{1-u^{|\alpha_{i}|}v^{\alpha_{i}\cdot S_{i}}}$ and for $2\leq i\leq t$, we have:
    $H_{\overset{\circ}{\sigma_{i}}}=\frac{u^{|\alpha_{i}|}v^{\alpha_{i}\cdot S_{i}}}{1-u^{|\alpha_{i}|}v^{\alpha_{i}\cdot S_{i}}}\cdot\frac{u^{|\alpha_{i-1}|}v^{\alpha_{i-1}\cdot S_{i-1}}}{1-u^{|\alpha_{i-1}|}v^{\alpha_{i-1}\cdot S_{i-1}}}$
    Also,
    $H_{\overset{\circ}{\sigma_{t+1}}}=\frac{1}{1-u^{|\alpha_{t}|}v^{\alpha_{t}\cdot S_{t+1}}}\sum_{i=1}^{p_{t}}u^{i+D_{i,t}+1}v^{i\cdot s_{t+1}}$
    where $SC(\alpha_{2,t}/\alpha_{1,t})=[[d_{1t},...,d_{p_{t},t}]]$, $D_{i,t}=\sum_{j=1}^{i}d_{j,t}$, $D_{0,t}=0$ and $S_{t+1}=(s_{t+1},0)$.
  • If $t>1$, set $T_{i}=\mathbf{R}_{\geq 0}\cdot\alpha_{i}$ with $\mathbf{R}_{\geq 0}(1,1)<T_{1}<...<T_{t}$ then
    $H=H_{T_{1}}+...+H_{T_{t}}+H_{\overset{\circ}{\sigma_{1}}}+...+H_{\overset{\circ}{\sigma_{t+1}}}$
    with $H_{T_{i}}=\frac{u^{|\alpha_{i}|}v^{\alpha_{i}\cdot S_{i}}}{1-u^{|\alpha_{i}|}v^{\alpha_{i}\cdot S_{i}}}$, and for $2\leq i\leq t+1$, $H_{\overset{\circ}{\sigma_{i}}}$ is as the previous point and
    $H_{\overset{\circ}{\sigma_{1}}}=\frac{1}{1-u^{|\alpha_{1}|}v^{\alpha_{1}\cdot S_{1}}}\sum_{i=1}^{p_{1}}\sum_{j=D_{i-1,1}+1}^{D_{i,1}}u^{i+j}v^{js_{1}}$
    where $S_{1}=(0,s_{1})$.
  • If $t>1$ with $T_{1}<\mathbf{R}_{\geq 0}(1,1)<T_{t}$ then
    $H=H_{T_{1}}+...+H_{T_{t}}+H_{\overset{\circ}{\sigma_{1}}}+...+H_{\overset{\circ}{\sigma_{t+1}}}$
    $H_{T_{i}}$ are as before, $H_{\overset{\circ}{\sigma_{i}}}$ as before for $2\leq i\leq t+1$ and finally,
    $H_{\overset{\circ}{\sigma_{1}}}=\frac{1}{1-u^{|\alpha_{1}|}v^{\alpha_{1}\cdot S_{1}}}\sum_{i=1}^{p_{1}}u^{i+D_{1,i}+1}v^{i\cdot s_{1}}$
    where $D_{1,i}=\sum_{j=1}^{i}d_{1,j}$, $SC(\alpha_{1,1}/\alpha_{2,1})=[[d_{11},...,d_{p_{1},1}]]$.
Proof.
This comes from the fact that $H$ has a decomposition:
$H=\sum_{i=1}^{t}H_{T_{i}}+\sum_{i=1}^{t+1}H_{\overset{\circ}{\sigma_{i}}}$
where $H_{T_{i}}$ is the sum over the lattice points who lie on $T_{i}\cap\mathbf{N}^{2}$ and $H_{\overset{\circ}{\sigma_{i}}}$ is the sum over the lattice points who lie in the interior of the cone $\sigma_{i}=T_{i-1}+T_{i}$. We have always:
$H_{T_{i}}=\sum_{s=1}^{\infty}u^{s|\alpha_{i}|}v^{s\alpha_{i}\cdot S_{i}}=\frac{u^{|\alpha_{i}|}v^{\alpha_{i}\cdot S_{i}}}{1-u^{|\alpha_{i}|}v^{\alpha_{i}\cdot S_{i}}}$
and for $2\leq i\leq t$,
$H_{\overset{\circ}{\sigma_{i}}}=\sum_{r,s\geq 1}u^{r|\alpha_{i}|+s|\alpha_{i-1}|}v^{(r\alpha_{i}+s\alpha_{i-1})\cdot S_{i}}=\frac{u^{|\alpha_{i}|}v^{\alpha_{i}\cdot S_{i}}}{1-u^{|\alpha_{i}|}v^{\alpha_{i}\cdot S_{i}}}\cdot\frac{u^{|\alpha_{i-1}|}v^{\alpha_{i-1}\cdot S_{i-1}}}{1-u^{|\alpha_{i-1}|}v^{\alpha_{i-1}\cdot S_{i-1}}}$
since $\alpha_{i-1}\cdot S_{i}=\alpha_{i-1}\cdot S_{i-1}$ for $\alpha_{i}\in\sigma_{i}$. For $H_{\overset{\circ}{\sigma_{1}}}$ and $H_{\overset{\circ}{\sigma_{t+1}}}$, there are multiple cases depending on whether $T(f_{1})\geq\mathbf{R}_{\geq 0}(1,1)$ or $T(f_{1})<\mathbf{R}_{\geq 0}(1,1)<T(f_{t})$.
  • if $T(f_{1})=\mathbf{R}_{\geq 0}(1,1)$ then $H_{\overset{\circ}{\sigma}_{1}}=0$ ;
  • if $T(f_{1})>\mathbf{R}_{\geq 0}(1,1)$ and $t=1$ then choose $\alpha=(p,q)$ the generator of the tropical variety of $f$ such that $gcd(p,q)=1$ $SC(q/p)=[[d_{1},...,d_{p}]]$. Also,
    $\overset{\circ}{\sigma}_{1}\cap B^{*}=\{\lambda+u:\lambda\in T(f)\cap\mathbf{N}^{2}_{0},u\in U\}$
    where $U=\{(1,1),...,(1,d_{1}),(2,d_{1}+1),...,(2,d_{1}+d_{2}),...,(p,d_{1}+...+d_{p-1}+1),...,(p,q-1)\}$ and we have
    $H_{\overset{\circ}{\sigma}_{1}}=\sum_{\lambda\in T(f)\cap\mathbf{N}_{0}^{2}}\>\ \sum_{\nu\in U}u^{|\lambda|+|\nu|}v^{[\lambda+\nu]\cdot S_{1}}$
    where $S_{1}=(0,s_{1})$ is the first vertex of the Newton polygon of $f$. As for $H_{\overset{\circ}{\sigma}_{2}}$ we have:
    $H_{\overset{\circ}{\sigma}_{2}}=\sum_{\lambda\in T(f)\cap\mathbf{N}_{0}^{2}}\>\ \sum_{\nu\in V}u^{|\lambda|+|\nu|}v^{[\lambda+\nu]\cdot S_{2}}$
    where $V=\{(1,d_{1}+1),(2,d_{1}+d_{2}+1),...,(p-1,d_{1}+...+d_{p-1}+1)\}$.
  • if $T(f_{1})>\mathbf{R}_{\geq 0}(1,1)$ and $t>1$ then $H_{\overset{\circ}{\sigma}_{1}}$ (resp. $H_{\overset{\circ}{\sigma}_{t+1}}$) is computed as $H_{\overset{\circ}{\sigma}_{1}}$ (resp. $H_{\overset{\circ}{\sigma}_{2}}$) in the previous point but with replacing $T(f)$ by $T(f_{1})$ (resp. by $T(f_{t})$) and $SC(q/p)$ by $SC(\alpha_{1,2}/\alpha_{1,1})$ (resp. $SC(\alpha_{t,2}/\alpha_{t,1})$).
  • if $T(f_{1})<\mathbf{R}_{\geq 0}(1,1)<T(f_{t})$, then $H_{\overset{\circ}{\sigma}_{t+1}}$ is computed as $H_{\overset{\circ}{\sigma}_{t+1}}$ in the previous point. As for $H_{\overset{\circ}{\sigma}_{1}}$ it is computed as such:
    $H_{\overset{\circ}{\sigma}_{1}}=\sum_{\lambda\in T(f)\cap\mathbf{N}_{0}^{2}}\>\ \sum_{u^{\prime}\in U^{\prime}}u^{|\lambda|+|u^{\prime}|}v^{[\lambda+u^{\prime}]\cdot S_{1}}$
    where $U^{\prime}$ is the same set as $U$ but where axis and ordinates of each lattice point are flipped.
When we compute each of $H_{\overset{\circ}{\sigma}_{1}}$ and $H_{\overset{\circ}{\sigma}_{t+1}}$ in detail we find the expression written in the statement of the theorem. The case where $T(f_{1})>T(f_{t})$ doesn’t exist because we have supposed
$f=f_{1}...f_{t}$
an oriented tropical decomposition and the case $T(f_{1})\leq T(f_{t})\leq\mathbf{R}_{\geq 0}(1,1)$ can be flipped to $\mathbf{R}_{\geq 0}(1,1)\leq T(g_{1})\leq T(g_{t})$ where $g_{i}(x,y)=f_{i}(y,x)$ for all $i$.
Finally, we compute the poles of the generating series $G$ of $f$ defining a Newton plane curve singularity.
Theorem 11.
Let $f$ define a Newton non-degenerate plane curve singularity. Denote by $G$ its generating series. Then $G=H+R$ (from the previous lemmas) is in $\mathbf{Z}(u,v)$ and its poles with respect to $(u,v)$ are:
$\{uv=1\}\cup\{u^{\frac{|\alpha_{l}|}{\Delta_{l}}}\cdot v^{\frac{\alpha_{l}}{\Delta_{l}}\cdot S_{l}}=e^{\frac{2ik\pi}{\Delta_{l}}}\}_{1\leq l\leq t\>\ 1\leq k\leq\Delta_{l}}$
where $\Delta_{l}=gcd(|\alpha_{l}|,\alpha_{l}\cdot S_{l})$.
Proof.
Recall $\alpha_{i}\cdot S_{i}=\nu_{\alpha_{i}}(f)$. $R=\frac{u}{1-uv}\sum_{i=1}^{t}r_{i}H_{T_{i}}$ where $H_{T_{i}}=\frac{u^{|\alpha_{i}|}v^{\nu_{\alpha_{i}}(f)}}{1-u^{|\alpha_{i}|}v^{\nu_{\alpha_{i}}(f)}}$. Moreover, $H=\sum_{i=1}^{t}H_{T_{i}}+\sum_{i=1}^{t+1}H_{\overset{\circ}{\sigma_{i}}}$ where for $2\leq i\leq t$, $H_{\overset{\circ}{\sigma_{i}}}=H_{T_{i}}H_{T_{i-1}}$ and for $j\in\{1,t+1\}$, write
$H_{\overset{\circ}{\sigma_{j}}}=\frac{P_{j}(u,v)}{(1-u^{|\alpha_{j}|}v^{\nu_{\alpha_{j}}(f)})}$
where $P_{j}\in\mathbf{C}[u,v]$. Thus, the set of poles $\mathcal{P}$ of $G$ verifies
$\mathcal{P}\subseteq\{uv=1\}\cup\{u^{\frac{|\alpha_{l}|}{\Delta_{l}}}\cdot v^{\frac{\alpha_{l}}{\Delta_{l}}\cdot S_{l}}=e^{\frac{2ik\pi}{\Delta_{l}}}\}_{1\leq l\leq t\>\ 1\leq k\leq\Delta_{l}}$
because we have the factorization into irreducible factors:
$u^{|\alpha_{l}|}v^{\alpha_{l}\cdot S_{l}}-1=\prod_{k=1}^{\Delta_{l}}(u^{\frac{|\alpha_{l}|}{\Delta_{l}}}\cdot v^{\frac{\alpha_{l}}{\Delta_{l}}\cdot S_{l}}-e^{\frac{2ik\pi}{\Delta_{l}}})$
  • If $T(f)=\mathbf{R}_{\geq 0}(1,1)$ with $f$ having two branches. Choose coordinates $(x,y)$ to have
    f=(yx+)(ycx+)f=(y-x+...)(y-cx+...)
    Then $s=2$ in
    $G=\frac{u^{2}v^{s}}{1-u^{2}v^{s}}+\frac{2u}{(1-uv)^{2}}\cdot\frac{u^{2}v^{2}}{1+uv}=-1+\frac{1/2}{1-uv}+\frac{1/2}{1+uv}+\frac{2u}{(1-uv)^{2}}\cdot\frac{u^{2}v^{2}}{1+uv}$
    has 1uv1-uv as a pole of multiplicity $2$, $1+uv$ of multiplicity $1$.
  • Otherwise, if also $(1,1)\notin supp(f)$, according to Observation 2, then $\nu_{\alpha}(f)>|\alpha|$ for all $\alpha\in T(f)$. Then,
    $G=H+R$
    with $H$ not having 1uv1-uv as a pole and $R$ having it with multiplicity $1$. Then $G$ has 1uv1-uv as a pole of multiplicity $1$. Also,
    $R=\frac{u}{1-uv}\sum_{i=1}^{t}r_{i}H_{T_{i}}$
    has $e^{\frac{2ik\pi}{\Delta_{l}}}-u^{\frac{|\alpha_{l}|}{\Delta_{l}}}\cdot v^{\frac{\alpha_{l}}{\Delta_{l}}\cdot S_{l}}$ as a pole of $R$ of multiplicity $1$. Since $G=H+R$ and $H$ doesn’t have 1uv1-uv as a pole then $G$ has $e^{\frac{2ik\pi}{\Delta_{l}}}-u^{\frac{|\alpha_{l}|}{\Delta_{l}}}\cdot v^{\frac{\alpha_{l}}{\Delta_{l}}\cdot S_{l}}$ as a pole.
  • Finally if $(1,1)\in supp(f)$ then the Newton Polygon of $f$ has two faces (ie $t=2$) with $(1,1)$ as a middle vertex. Moreover, $f$ has two branches since $(1,1)\in supp(f)$ and for all $\alpha\in T(f)$, $\nu_{\alpha}(f)=|\alpha|$. So $\alpha$ and $S=(1,1)$ should correspond to poles of the form
    $\{uv=e^{\frac{2ik\pi}{|\alpha|}}\}_{1\leq k\leq|\alpha|}$
    $T(f)=\mathbf{R}_{\geq 0}\cdot\alpha\cup\mathbf{R}_{\geq 0}\cdot\beta$ where $\alpha=(a,1)$, $\beta=(1,b)$ since $(1,1)\in supp(f)$.
    $R=(1+\frac{u}{1-uv})(\frac{u^{|\alpha|}v^{|\alpha|}}{1-u^{|\alpha|}v^{|\alpha|}}+\frac{u^{|\beta|}v^{|\beta|}}{1-u^{|\beta|}v^{|\beta|}})$
    and according to Lemma 7,
    $H=\frac{(uv)^{|\alpha|+|\beta|}}{(1-u^{|\alpha|}v^{|\alpha|})(1-u^{|\beta|}v^{|\beta|})}+\frac{u^{|\alpha|}v^{|\alpha|}+u^{1+|\alpha|}v^{s_{1}}}{1-u^{|\alpha|}v^{|\alpha|}}+\frac{u^{|\beta|}v^{|\beta|}+u^{1+|\beta|}v^{s_{3}}}{1-u^{|\beta|}v^{|\beta|}}$
    where $S_{1}=(0,s_{1})$ and $S_{3}=(s_{3},0)$ are the first and last vertex of the Newton polygon of $f$. This shows that the poles of $G$ are
    $\{uv=e^{\frac{2ik\pi}{|\alpha|}}\}_{1\leq k\leq|\alpha|}\cup\{uv=e^{\frac{2ik\pi}{|\beta|}}\}_{1\leq k\leq|\beta|}$
This concludes the proof. ∎